### v1

# Quantum Entanglement Geometry on Severi–Brauer Schemes: Subsystem Reductions of Azumaya Algebras

Abstract.
We formulate pure-state entanglement in families as a geometric obstruction. In standard quantum information, entanglement is defined relative to a chosen tensor-product factorization of a fixed Hilbert space. In contrast, for a twisted family of pure-state spaces, which can be described by Azumaya algebras $\mathcal{A}$ of degree $n$ on $X$ and their Severi–Brauer schemes
$SB(\mathcal{A})=P\times^{\operatorname{PGL}_{n}}\mathbb{P}^{n-1}\to X,$
such a subsystem choice may fail to globalize. We formalize this algebro-geometrically: fixing a factorization type $\mathbf{d}=(d_{1},\dots,d_{s})$ with $n=\prod_{i}d_{i}$, the existence of a global product-state locus of type $\mathbf{d}$ is equivalent to a reduction of the underlying $\operatorname{PGL}_{n}$–torsor $P\to X$ to the stabilizer $G_{\mathbf{d}}\subset\operatorname{PGL}_{n}$. Thus, entanglement is the obstruction to the existence of a relative Segre subscheme inside $SB(\mathcal{A})$.
Writing $\Sigma_{\mathbf{d}}\subset\mathbb{P}^{n-1}$ for the Segre variety, we call a reduction to $G_{\mathbf{d}}$ a $\mathbf{d}$–subsystem structure. Our first main result identifies the moduli of $\mathbf{d}$–subsystem structures with the quotient $P/G_{\mathbf{d}}$. Moreover, we realize naturally $P/G_{\mathbf{d}}$ as a locally closed subscheme of the relative Hilbert scheme,
$\operatorname{Hilb}^{\Sigma_{\mathbf{d}}}\!\bigl(SB(\mathcal{A})/X\bigr)\ \subset\ \operatorname{Hilb}\bigl(SB(\mathcal{A})/X\bigr),$
parametrizing relative closed subschemes fppf-locally isomorphic to $\Sigma_{\mathbf{d}}\times X$.
In the bipartite case $\mathbf{d}=(d_{A},d_{B})$, $G_{\mathbf{d}}$–invariant determinantal geometry globalizes the Schmidt-rank filtration, with base-change compatible resolutions and numerical invariants that are constant on fibers. Finally, we study Brauer-theoretic constraints on the existence of subsystem structures, including an $\mathrm{lcm}(\mathbf{d})$–torsion obstruction and examples showing that $\mathbf{d}$–reducibility depends on the underlying $\operatorname{PGL}_{n}$–torsor, not only on the Brauer class.
2020 Mathematics Subject Classification:
Primary 14F22; Secondary 16K50, 14C05, 14M12, 14M15, 14L30, 81P40.
The author expresses gratitude to Steven Rayan for his careful reading of the manuscript and for his invaluable comments. This work was partially supported by the NSF under Grant No. OSI-2328774.
1. Introduction
Motivation: Entanglement as geometry
Entanglement is a defining feature of genuinely quantum behavior. For pure states, it can be expressed in algebro–geometric terms once a tensor-product background is fixed: for a factorization
$\mathcal{H}\cong\mathcal{H}_{1}\otimes\cdots\otimes\mathcal{H}_{s},\qquad\dim(\mathcal{H}_{i})=d_{i},\ \ \prod_{i}d_{i}=n,$
the projective space $\mathbb{P}(\mathcal{H})\cong\mathbb{P}^{n-1}$ contains the Segre variety
$\Sigma_{\mathbf{d}}:=\mathbb{P}^{d_{1}-1}\times\cdots\times\mathbb{P}^{d_{s}-1}\hookrightarrow\mathbb{P}^{n-1},$
which is precisely the locus of product states of type $\mathbf{d}=(d_{1},\dots,d_{s})$. In the bipartite case, the determinantal rank loci recover the Schmidt-rank stratification and provide a geometric entanglement filtration.
The starting point of this paper is that the usual definition of entanglement presupposes the existence of a global subsystem structure. In many geometric and physical situations one encounters families of pure-state spaces over a base $X$, where the fibers are projective spaces only locally and may be globally twisted. In such a family, one can often choose a tensor factorization on a simply connected chart, but it is not automatic that these local subsystem choices transport and glue globally. When they do not, there is no globally defined product-state locus, hence no globally meaningful notion of “entangled vs. product” states for the family.
Known facts: Severi–Brauer families as twisted pure-state spaces
Let $k$ be an algebraically closed field of characteristic 0, and let $X$ be a smooth quasi-projective kk–scheme. A degree-$n$ Azumaya algebra $\mathcal{A}$ determines a principal $\operatorname{PGL}_{n}$–torsor $P\to X$ and the associated Severi–Brauer scheme
$SB(\mathcal{A})=P\times^{\operatorname{PGL}_{n}}\mathbb{P}^{n-1}\to X,$
which is fppf-locally a trivial $\mathbb{P}^{n-1}$–bundle. From the quantum-information viewpoint, $SB(\mathcal{A})\to X$ is a natural algebro–geometric model for a family of pure-state spaces whose projective trivializations may carry nontrivial twisting.
Grothendieck’s theory of Azumaya algebras and the Brauer group provides the natural language for such families [Grothendieck1968BrauerI, Grothendieck1968BrauerII, Grothendieck1968BrauerIII] (see also [Azumaya1951MaximallyCentral, AuslanderGoldman1960Brauer]). These tools successfully describe the background twisting of the projective state bundle. However, they do not by themselves supply a globally coherent notion of quantum entanglement, and thus they do not single out a canonical Segre locus inside $SB(\mathcal{A})$.
The remaining problems.
Fix a factorization type $\mathbf{d}=(d_{1},\dots,d_{s})$ with $n=\prod_{i}d_{i}$, let $\Sigma_{\mathbf{d}}\subset\mathbb{P}^{n-1}$ be the Segre variety, and denote by $G_{\mathbf{d}}\subset\mathrm{PGL}_{n}$ its stabilizer (Definition 2.3).
Question.
Our central questions addressed in this paper are summarized as follows:
  1. (1)
    How to address/define quantum entanglement in a Severi–Brauer fibration $\pi:SB(\mathcal{A})\to X$?
  2. (2)
    When does $SB(A)$ contain a relative Segre subscheme of type $\mathbf{d}$ (when does the associated $\mathrm{PGL}_{n}$-torsor admit a reduction to $G_{\mathbf{d}}$)? In terms of quantum information, when does $SB(\mathcal{A})$ admits a family of globally compatible product states?
  3. (3)
    When it does not, what obstructs such a globalization? What are consistent views of such obstructions in quantum information theory?
Unlike the standard (untwisted) setting in which a tensor-product background is fixed from the outset, here the subsystem choice itself is a genuine local-to-global datum: although $\Sigma_{\mathbf{d}}$ is visible after any local trivialization $SB(A)|_{U}\simeq\mathbb{P}^{n-1}\times U$, the twisting of the family may prevent these local product-state loci from gluing. This issue is structurally significant. In the absence of a globally well-defined product-state locus, it is not possible to formulate – uniformly over $X$ – quantum entanglement, nor to develop its associated relative geometric framework.
Our approaches
(1) Stabilizer reductions and entanglement as an obstruction class.
Our first main result identifies this reduction problem with the geometric existence of a descended Segre locus.
We prove that the above question is governed purely by torsor reduction: a relative Segre subscheme of type $d$ exists in $SB(A)$ if and only if the $\operatorname{PGL}_{n}$–torsor $P\to X$ admits a reduction of structure group to $G_{\mathbf{d}}$ (Corollary 2.13). We call such a reduction a dd–subsystem structure (Definition 2.6). In this sense, entanglement in families is exactly the obstruction to reducing $P$ to $G_{d}$.
In the language of quantum information, entanglement in families is the obstruction to globalizing product states. When the reduction fails, transition functions necessarily exit the stabilizer $G_{\mathbf{d}}$ and the local product-state loci cannot be transported globally. In particular, a state that is locally a product state can become entangled after gluing along the base.
The obstruction is naturally tied to Brauer-theoretic data. We introduce $\operatorname{Br}_{\mathbf{d}}(X)\subset\operatorname{Br}(X)$, the subset of Brauer classes that admit at least one degree-$n$ Azumaya representative equipped with a $\mathbf{d}$–subsystem structure (Definition 2.7). We prove a general torsion constraint
[…]
and develop examples showing that the obstruction can be genuinely finer than torsion, and can depend on the underlying $\operatorname{PGL}_{n}$–torsor rather than only on the Brauer class.
(2) A moduli space of subsystem choices as a Hilbert–scheme locus
Subsystem structures are not merely a yes/no property: when they exist, they form a natural moduli space. We show that the functor of $G_{\mathbf{d}}$–reductions of $P$ is represented by the quotient $P/G_{\mathbf{d}}$, which embeds naturally as a locally closed subscheme of the relative Hilbert scheme:
[…]
In characteristic 0 this locus is smooth over $X$ and étale-locally a trivial $\operatorname{PGL}_{n}/G_{\mathbf{d}}$–fibration (Proposition 2.20). For $\mathbf{d}=(2,2)$ the relevant Hilbert component is $\mathbb{P}^{9}$, and the boundary of the open Segre orbit is the (discriminant) divisor of singular quadrics (Proposition 2.23 and Corollary 2.24). This Hilbert-scheme viewpoint supplies a canonical compactification and makes degeneration phenomena intrinsic to the Severi–Brauer background.
(3) From reductions to entanglement geometry
From Section 3 onward we specialize to the bipartite case $\mathbf{d}=(d_{A},d_{B})$. A key point is that determinantal entanglement geometry is not available on a general Severi–Brauer family without extra structure: a twisted $\mathbb{P}^{n-1}$–bundle does not come with a global tensor identification. Once a bipartite subsystem structure is fixed, we show that the classical rank loci globalize to a flat filtration, which we call entanglement filtration inspired from quantum information,
$\Sigma_{\leq 1}(\mathcal{A},\mathbf{d})\subset\Sigma_{\leq 2}(\mathcal{A},\mathbf{d})\subset\cdots\subset SB(\mathcal{A}),$
with base-change compatible incidence resolutions and a uniform singularity theory (Theorem 3.10). Moreover, the resulting intersection-theoretic and numerical data are fiberwise constant and computable (Theorem 5.5, Corollary 3.13). In this way, the background geometry controls which entanglement structures exist globally, and when they do exist, the family carries universal invariants that are stable under base change.
(4) Explicit obstructions and new background-sensitive entanglement signatures
We complement the general theory with concrete obstruction computations and examples. For instance, on $X=\mathbb{G}_{\mathrm{m}}^{2}$ the generic Kummer $p^{2}$–symbol class does not admit a global $(p,p)$ subsystem structure (Theorem 4.7), while we also construct nontrivial twisted degree-$p^{2}$ families that do admit a global $(p,p)$ reduction (Proposition 4.16 and Example 3.14). We further treat situations where different local tensor types occur on different charts (Example 4.14), so that the effective “particle number” depends on the locally chosen subsystem.
A recurrent theme is that subsystem reducibility is not a Brauer invariant: even for $\beta=0$ it can depend on the finer $\operatorname{PGL}_{n}$–torsor class (Proposition 4.27), and on $\mathbb{P}^{1}$ we obtain a complete splitting-type classification in the bipartite case (Theorem 4.30).
(5) Quantum-information interpretation: entangling monodromy.
When $k=\mathbb{C}$, local trivializations on $X^{\mathrm{an}}$ can acquire monodromy in $\operatorname{PGL}_{n}(\mathbb{C})$. Failure of reduction to $G_{\mathbf{d}}$ forces some transition elements outside $G_{\mathbf{d}}$, producing an entangling monodromy acting on locally defined product states (Proposition 4.11). In a concrete 44–level model this holonomy is locally equivalent to a $\mathrm{CNOT}$-type gate (Theorem 4.10). Appendix A gives a spin-chain toy model on a torus where a locally product ground state becomes entangled after gluing, realizing the obstruction mechanism in a physical setting.
We distinguish two local-to-global obstruction problems:
  • (I)
    Obstruction of structure (this paper). Whether a tensor-product type $\mathbf{d}$ can be chosen globally on the Severi–Brauer family, detected by the existence of a reduction of $P$ to $G_{\mathbf{d}}$ (equivalently, by the existence of a descended Segre subfibration).
  • (II)
    Obstruction of states (e.g. [Ikeda:2025mgj, Abramsky:2011sbx]). After fixing the background and a subsystem structure, whether local state data glue to a global state, and whether that global state is unique.
Both are local-to-global in nature, but they occur at different layers (structure versus states).
Organization
Section 2 introduces subsystem structures as $G_{\mathbf{d}}$–reductions and the subset $\operatorname{Br}_{\mathbf{d}}(X)\subset\operatorname{Br}(X)$. Section 2.3 identifies the moduli of reductions with a canonically defined Hilbert–locus (Theorem 2.21). In the bipartite case, Section 3 develops the relative determinantal filtration, including uniform resolutions and numerical invariants. Examples and obstructions appear in Section 4, and some numerical invariants are collected in Section 5.
2. Azumaya algebras and subsystem structures
2.1. Azumaya algebras and subsystem structures
Definition 2.1 (Azumaya algebra of degree $n$).
An Azumaya algebra of degree $n$ on $X$ is a sheaf of $\mathcal{O}_{X}$–algebras $\mathcal{A}$ which is fppf-locally isomorphic to $M_{n}(\mathcal{O}_{X})$. It is associated to a principal $\operatorname{PGL}_{n}$–torsor $P\to X$ via
$\mathcal{A}\;\cong\;P\times^{\operatorname{PGL}_{n}}\bigl(M_{n}(k)\otimes_{k}\mathcal{O}_{X}\bigr)\;\cong\;P\times^{\operatorname{PGL}_{n}}M_{n}(\mathcal{O}_{X}),$
where $\operatorname{PGL}_{n}$ acts on $M_{n}(k)$ by conjugation and acts trivially on $\mathcal{O}_{X}$ (so the action on $M_{n}(k)\otimes_{k}\mathcal{O}_{X}$ is $\mathcal{O}_{X}$–linear).
Definition 2.2 (Severi–Brauer scheme).
Let $\mathcal{A}$ be an Azumaya algebra of degree $n$ on $X$. Its Severi–Brauer scheme is the XX–scheme
$\pi:SB(\mathcal{A})\longrightarrow X$
characterized fppf-locally by $SB(\mathcal{A})|_{U}\cong\mathbb{P}^{n-1}\times U$ whenever $\mathcal{A}|_{U}\cong M_{n}(\mathcal{O}_{U})$. It represents locally free right ideals $I\subset\mathcal{A}$ of reduced rank $1$.
We write $\operatorname{Br}(X):=H^{2}_{\mathrm{fppf}}(X,\mathbb{G}_{m})_{\mathrm{tors}}$ for the Brauer group and denote by $\beta=[\mathcal{A}]\in\operatorname{Br}(X)$ the Brauer class of $\mathcal{A}$.
Let $\Sigma_{\mathbf{d}}\subset\mathbb{P}^{n-1}$ denote the Segre variety ($\mathbf{d}$ product-state locus)
$\Sigma_{\mathbf{d}}:=\mathbb{P}^{d_{1}-1}\times\cdots\times\mathbb{P}^{d_{s}-1}\hookrightarrow\mathbb{P}(\,k^{d_{1}}\otimes\cdots\otimes k^{d_{s}}\,)\cong\mathbb{P}^{n-1}.$
Definition 2.3 (Stabilizer).
Define $G_{\mathbf{d}}\subset\operatorname{PGL}_{n}$ to be the (scheme-theoretic) stabilizer of the Segre subscheme:
$G_{\mathbf{d}}:=\{\,g\in\operatorname{PGL}_{n}\mid g(\Sigma_{\mathbf{d}})=\Sigma_{\mathbf{d}}\,\}.$
Remark 2.4.
The connected component of $G_{\mathbf{d}}$ contains the image of the tensor product representation $\prod_{i}\operatorname{GL}_{d_{i}}\to\operatorname{GL}_{n}\to\operatorname{PGL}_{n}$. If some factors have equal dimension, $G_{\mathbf{d}}$ may contain additional discrete symmetries permuting the equal Segre factors. None of our results depend on choosing an ordering convention.
Remark 2.5 (Structure of the stabilizer).
It is classical that the (projective) linear automorphisms preserving the product-state locus $\Sigma_{\mathbf{d}}\subset\mathbb{P}^{n-1}$ form $\bigl(\prod_{i}\operatorname{PGL}_{d_{i}}\bigr)\rtimes\Pi$, where $\Pi$ permutes the equal-dimensional factors. In particular, $G_{\mathbf{d}}$ is smooth in $\mathrm{char}(k)=0$. See e.g. Westwick [westwick1967transformations] (and a modern summary in [2024arXiv240716767G, Prop. 2.2]).
Definition 2.6 (Subsystem structure of type $\mathbf{d}$).
Let $\mathcal{A}$ be an Azumaya algebra of degree $n$ with associated $\operatorname{PGL}_{n}$–torsor $P\to X$ ($\mathbb{P}^{n-1}$–bundle). A subsystem structure of type $\mathbf{d}$ on $\mathcal{A}$ is a reduction of structure group of $P$ to $G_{\mathbf{d}}$, consisting of a principal $G_{\mathbf{d}}$–torsor $P_{\mathbf{d}}\to X$ and an identification
$P\;\cong\;P_{\mathbf{d}}\times^{G_{\mathbf{d}}}\operatorname{PGL}_{n}.$
Definition 2.7 ($\mathbf{d}$–admissible Brauer classes).
Let $\mathbf{d}=(d_{1},\dots,d_{s})$ with $n=\prod_{i}d_{i}$ and let $G_{\mathbf{d}}\subset\operatorname{PGL}_{n}$ be the stabilizer. Define
$\operatorname{Br}_{\mathbf{d}}(X):=\mathrm{Im}\Bigl(H^{1}(X,G_{\mathbf{d}})\to H^{1}(X,\operatorname{PGL}_{n})\xrightarrow{\delta}\operatorname{Br}(X)\Bigr),$
where $\delta$ is the connecting morphism associated to $1\to\mathbb{G}_{m}\to\operatorname{GL}_{n}\to\operatorname{PGL}_{n}\to 1$ in fppf cohomology.
Here $\beta\in\operatorname{Br}_{\mathbf{d}}(X)$ means that there exists at least one $\operatorname{PGL}_{n}$–torsor (Azumaya algebra of degree $n$) representing $\beta$ that admits a $G_{\mathbf{d}}$–reduction.
Definition 2.8 (Brauer-theoretic obstruction to $\mathbf{d}$–subsystem structures).
A Brauer class $\beta\in\operatorname{Br}(X)$ is $\mathbf{d}$–inadmissible if $\beta\notin\operatorname{Br}_{\mathbf{d}}(X)$, meaning that no Azumaya algebra of degree $n$ representing $\beta$ admits a $\mathbf{d}$–subsystem structure.
Remark 2.9.
In the quantum-information interpretation, $\mathbf{d}$–inadmissibility means that the twisting encoded by $\beta$ forbids the existence of a global notion of “product states of type $\mathbf{d}$” on the family $SB(\mathcal{A})\to X$. Thus such a state is intrinsically entangled. See Example 4.14.
Theorem 2.10.
Let $\mathcal{A}$ be an Azumaya algebra of degree $n$ with associated $\operatorname{PGL}_{n}$–torsor $P\to X$. Fix $\mathbf{d}$ and $G_{\mathbf{d}}\subset\operatorname{PGL}_{n}$.
  1. (i)
    Torsor level: The algebra $\mathcal{A}$ admits a $\mathbf{d}$–subsystem structure if and only if
    $P]\in\mathrm{Im}\bigl(H^{1}(X,G_{\mathbf{d}})\to H^{1}(X,\operatorname{PGL}_{n})\bigr).$
  2. (ii)
    Brauer level: If $\mathcal{A}$ admits a $\mathbf{d}$–subsystem structure, then its Brauer class $\mathcal{A}]=\delta([P])$ lies in $\operatorname{Br}_{\mathbf{d}}(X)$. Conversely, $\beta\in\operatorname{Br}_{\mathbf{d}}(X)$ guarantees the existence of some degree-$n$ Azumaya algebra $\mathcal{A}^{\prime}$ representing $\beta$ that admits a $\mathbf{d}$–subsystem structure, but this need not hold for every representative of $\beta$.
Proof.
This is the standard nonabelian cohomological criterion for reductions of structure group, applied to the inclusion $G_{\mathbf{d}}\hookrightarrow\operatorname{PGL}_{n}$. The Brauer-theoretic statements follow by applying the boundary map $\delta$. ∎
2.2. From subsystem structures to a descended product-state locus
Definition 2.11.
Assume $\mathcal{A}$ is equipped with a subsystem structure $P_{\mathbf{d}}\to X$ (Definition 2.6). Define the associated XX–scheme
$\Sigma_{\mathbf{d}}(\mathcal{A})\;:=\;P_{\mathbf{d}}\times^{G_{\mathbf{d}}}\Sigma_{\mathbf{d}}\;\longrightarrow\;X.$
By construction, $\Sigma_{\mathbf{d}}(\mathcal{A})$ is fppf-locally isomorphic to $\Sigma_{\mathbf{d}}\times U\to U$.
Proposition 2.12 (Descended product-state embedding).
There is a natural closed immersion over $X$
$\iota_{\mathbf{d}}:\Sigma_{\mathbf{d}}(\mathcal{A})\hookrightarrow SB(\mathcal{A})$
which is fppf-locally identified with the classical Segre embedding $\Sigma_{\mathbf{d}}\hookrightarrow\mathbb{P}^{n-1}$.
Proof.
Choose an fppf cover $U\to X$ trivialising the $G_{\mathbf{d}}$–torsor $P_{\mathbf{d}}$. After choosing a trivialization of $P_{\mathbf{d}}|_{U}$, we obtain isomorphisms
$SB(\mathcal{A})|_{U}\cong\mathbb{P}^{n-1}\times U,\qquad\Sigma_{\mathbf{d}}(\mathcal{A})|_{U}\cong\Sigma_{\mathbf{d}}\times U,$
and hence the classical Segre embedding gives a closed immersion on $U$.
On overlaps $U\times_{X}U$, the two trivializations differ by a transition function with values in $G_{\mathbf{d}}$, and by definition $G_{\mathbf{d}}$ preserves the Segre subscheme $\Sigma_{\mathbf{d}}\subset\mathbb{P}^{n-1}$. Therefore the local Segre maps are compatible on overlaps and glue to a global closed immersion $\Sigma_{\mathbf{d}}(\mathcal{A})\hookrightarrow SB(\mathcal{A})$. ∎
Corollary 2.13 (Subsystem structures $\Longleftrightarrow$ relative Segre subschemes).
Let $\mathcal{A}$ be an Azumaya algebra of degree $n$ on $X$, with Severi–Brauer scheme $\pi:SB(\mathcal{A})\to X$. Fix a factorization type $\mathbf{d}$ and the Segre subscheme $\Sigma_{\mathbf{d}}\subset\mathbb{P}^{n-1}$.
A $\mathbf{d}$–subsystem structure on $\mathcal{A}$ exists if and only if there exists a closed subscheme
$\Sigma\ \hookrightarrow\ SB(\mathcal{A})$
flat over $X$ such that fppf-locally on $X$ the pair $(SB(\mathcal{A}),\Sigma)$ is isomorphic to $(\mathbb{P}^{n-1}\times X,\ \Sigma_{\mathbf{d}}\times X)$, with $\Sigma$ identified with the standard Segre subscheme.
Moreover, when such $\Sigma$ exists it determines (up to unique isomorphism) a unique $G_{\mathbf{d}}$–reduction $P_{\mathbf{d}}\to X$ of the associated $\operatorname{PGL}_{n}$–torsor $P\to X$, and the induced subscheme $\Sigma_{\mathbf{d}}(\mathcal{A})=P_{\mathbf{d}}\times^{G_{\mathbf{d}}}\Sigma_{\mathbf{d}}$ of Definition 2.11 identifies canonically with $\Sigma$.
Proof.
The forward implication is Proposition 2.12.
Conversely, choose an fppf cover $U\to X$ trivializing $P$, so that $SB(\mathcal{A})|_{U}\simeq\mathbb{P}^{n-1}\times U$. By assumption, after refining $U$ we may identify $\Sigma|_{U}$ with $\Sigma_{\mathbf{d}}\times U$. On overlaps $U\times_{X}U$, the transition functions of $SB(\mathcal{A})$ are $\operatorname{PGL}_{n}$–valued maps $g_{ij}$, and the fact that the Segre subschemes glue forces $g_{ij}$ to preserve $\Sigma_{\mathbf{d}}$. Hence $g_{ij}$ takes values in $G_{\mathbf{d}}$, giving a $G_{\mathbf{d}}$–valued cocycle and therefore a $G_{\mathbf{d}}$–reduction $P_{\mathbf{d}}\to X$. By construction, the associated subscheme $P_{\mathbf{d}}\times^{G_{\mathbf{d}}}\Sigma_{\mathbf{d}}$ recovers $\Sigma$. ∎
2.3. Subsystem-structure loci
2.3.1. Definitions
Fix an Azumaya algebra $\mathcal{A}$ of degree $n$ on $X$, and let $P\to X$ be the associated $\operatorname{PGL}_{n}$–torsor so that $SB(\mathcal{A})\cong P\times^{\operatorname{PGL}_{n}}\mathbb{P}^{n-1}$. Fix a subsystem (factorization) type $\mathbf{d}=(d_{1},\dots,d_{s})$ with $n=\prod_{i}d_{i}$ and the stabilizer $G_{\mathbf{d}}\subset\operatorname{PGL}_{n}$ (Definition 2.3).
Since $X$ is noetherian and $\pi:SB(\mathcal{A})\to X$ is projective, the full Hilbert functor
$\begin{aligned} \operatorname{Hilb}\bigl(SB(\mathcal{A})/X\bigr)&:(\mathbf{Sch}/X)^{\mathrm{op}}\to\mathbf{Sets},\\ (T\xrightarrow{f}X)&\longmapsto\left\{\Sigma_{T}\subset SB(\mathcal{A})\times_{X}T\;\middle|\;\Sigma_{T}\text{ is a closed subscheme flat over }T\right\}\end{aligned}$
is representable by a scheme, denoted by $\operatorname{Hilb}\bigl(SB(\mathcal{A})/X\bigr)$.
Fix a relatively ample invertible sheaf $\mathcal{L}$ on $SB(\mathcal{A})$. For each Hilbert polynomial $\mathsf{P}$ (with respect to $\mathcal{L}$), let $\operatorname{Hilb}^{\mathsf{P}}_{\mathcal{L}}\!\bigl(SB(\mathcal{A})/X\bigr)$ denote the corresponding (projective) Hilbert scheme. Then the full Hilbert scheme decomposes as a disjoint union of open-and-closed subschemes
$\operatorname{Hilb}\bigl(SB(\mathcal{A})/X\bigr)\;=\;\coprod_{\mathsf{P}}\operatorname{Hilb}^{\mathsf{P}}_{\mathcal{L}}\bigl(SB(\mathcal{A})/X\bigr).$
We will only use the existence of the full Hilbert scheme and the fact that each $\operatorname{Hilb}^{\mathsf{P}}_{\mathcal{L}}$ is projective over $X$.
Definition 2.14 (Reduction functor (subsystem structures in families)).
Let $\mathbf{Sch}/X$ be the category of schemes over $X$. Define a functor
$\mathfrak{Red}_{\mathbf{d}}(\mathcal{A}):(\mathbf{Sch}/X)^{\mathrm{op}}\to\mathbf{Sets}$
by sending $f:T\to X$ to the set of $G_{\mathbf{d}}$–reductions of $P_{T}:=P\times_{X}T$, namely pairs
$(P_{\mathbf{d},T},\ \alpha_{T})$
where $P_{\mathbf{d},T}\to T$ is a principal $G_{\mathbf{d}}$–torsor and $\alpha_{T}:P_{\mathbf{d},T}\times^{G_{\mathbf{d}}}\operatorname{PGL}_{n}\xrightarrow{\sim}P_{T}$ is an isomorphism of $\operatorname{PGL}_{n}$–torsors, modulo the usual notion of isomorphism of reductions.
Definition 2.15.
Define a subfunctor
$\mathfrak{Hilb}^{\Sigma_{\mathbf{d}}}_{\mathcal{A}}\ \subset\ \operatorname{Hilb}\bigl(SB(\mathcal{A})/X\bigr)$
as follows. For $f:T\to X$, $\mathfrak{Hilb}^{\Sigma_{\mathbf{d}}}_{\mathcal{A}}(T)$ consists of closed subschemes
$\Sigma_{T}\ \subset\ SB(\mathcal{A})\times_{X}T$
flat over $T$, such that there exists an fppf cover $\{T_{i}\to T\}$ with:
  1. (i)
    $SB(\mathcal{A})\times_{X}T_{i}\ \cong\ \mathbb{P}^{n-1}\times T_{i}$ over $T_{i}$ ($\mathcal{A}_{T_{i}}\cong M_{n}(\mathcal{O}_{T_{i}})$, so the induced $\operatorname{PGL}_{n}$–torsor on $T_{i}$ is trivial), and
  2. (ii)
    under such an identification, $\Sigma_{T}|_{T_{i}}$ is identified with $\Sigma_{\mathbf{d}}\times T_{i}\subset\mathbb{P}^{n-1}\times T_{i}$.
2.3.2. The Segre orbit inside the full Hilbert scheme of $\mathbb{P}^{n-1}$
Let $\operatorname{Hilb}(\mathbb{P}^{n-1})$ denote the full Hilbert scheme of $\mathbb{P}^{n-1}$ (the disjoint union of $\operatorname{Hilb}^{\mathsf{P}}(\mathbb{P}^{n-1})$ over all Hilbert polynomials $\mathsf{P}$} with respect to $\mathcal{O}_{\mathbb{P}^{n-1}}(1)$). The natural action of $\operatorname{PGL}_{n}$ on $\mathbb{P}^{n-1}$ induces an action on $\operatorname{Hilb}(\mathbb{P}^{n-1})$.
Let $\Sigma_{\mathbf{d}}]\in\operatorname{Hilb}(\mathbb{P}^{n-1})(k)$ be the point corresponding to the Segre subscheme, and let $\mathscr{O}_{\Sigma_{\mathbf{d}}}\subset\operatorname{Hilb}(\mathbb{P}^{n-1})$ denote its $\operatorname{PGL}_{n}$–orbit. Note that $\mathscr{O}_{\Sigma_{\mathbf{d}}}$ is automatically contained in the open-and-closed component of $\operatorname{Hilb}(\mathbb{P}^{n-1})$ containing $\Sigma_{\mathbf{d}}]$.
Lemma 2.16 (Segre orbit as a homogeneous space).
Let $\mathscr{O}_{\Sigma_{\mathbf{d}}}\subset\operatorname{Hilb}(\mathbb{P}^{n-1})$ be the $\operatorname{PGL}_{n}$–orbit of $\Sigma_{\mathbf{d}}]$. Then:
  1. (i)
    The stabilizer of $\Sigma_{\mathbf{d}}]$ in $\operatorname{PGL}_{n}$ is precisely $G_{\mathbf{d}}$.
  2. (ii)
    The orbit morphism
    $\phi:\ \operatorname{PGL}_{n}\longrightarrow\operatorname{Hilb}(\mathbb{P}^{n-1}),\qquad g\longmapsto g\cdot[\Sigma_{\mathbf{d}}]$
    factors through the quotient to give a morphism
    $\bar{\phi}:\ \operatorname{PGL}_{n}/G_{\mathbf{d}}\ \longrightarrow\ \operatorname{Hilb}(\mathbb{P}^{n-1})$
    which is a locally closed immersion with image $\mathscr{O}_{\Sigma_{\mathbf{d}}}$. In particular, $\mathscr{O}_{\Sigma_{\mathbf{d}}}$ is a smooth locally closed subscheme and
    $\mathscr{O}_{\Sigma_{\mathbf{d}}}\ \cong\ \operatorname{PGL}_{n}/G_{\mathbf{d}}.$
Proof.
(i) This is immediate from the definition of $G_{\mathbf{d}}$ as the scheme-theoretic stabilizer of the Segre subscheme $\Sigma_{\mathbf{d}}\subset\mathbb{P}^{n-1}$ (Definition 2.3).
(ii) Since $G_{\mathbf{d}}$ stabilizes $\Sigma_{\mathbf{d}}]$, the orbit map $\phi$ is constant on right cosets of $G_{\mathbf{d}}$, hence factors through a unique morphism $\bar{\phi}:\operatorname{PGL}_{n}/G_{\mathbf{d}}\to\operatorname{Hilb}(\mathbb{P}^{n-1})$.
We claim that $\bar{\phi}$ is a monomorphism. Let $T$ be a scheme and let $x_{1},x_{2}\in(\operatorname{PGL}_{n}/G_{\mathbf{d}})(T)$ with $\bar{\phi}(x_{1})=\bar{\phi}(x_{2})$. After an fppf cover $T^{\prime}\to T$ we may choose lifts $g_{1},g_{2}\in\operatorname{PGL}_{n}(T^{\prime})$ mapping to $x_{1}|_{T^{\prime}}$ and $x_{2}|_{T^{\prime}}$. The equality $\bar{\phi}(x_{1})=\bar{\phi}(x_{2})$ implies
$g_{1}(\Sigma_{\mathbf{d}}\times T^{\prime})\;=\;g_{2}(\Sigma_{\mathbf{d}}\times T^{\prime})\quad\text{as closed subschemes of }\mathbb{P}^{n-1}\times T^{\prime},$
hence $g_{2}^{-1}g_{1}$ stabilizes $\Sigma_{\mathbf{d}}\times T^{\prime}$, hence $g_{2}^{-1}g_{1}\in G_{\mathbf{d}}(T^{\prime})$ by (i). Therefore $g_{1}$ and $g_{2}$ define the same $T^{\prime}$–point of $\operatorname{PGL}_{n}/G_{\mathbf{d}}$, so $x_{1}|_{T^{\prime}}=x_{2}|_{T^{\prime}}$, and hence $x_{1}=x_{2}$. Thus $\bar{\phi}$ is a monomorphism.
A standard result implies that a monomorphism locally of finite type is an immersion. Hence $\bar{\phi}$ is an immersion. Its image is exactly the orbit $\mathscr{O}_{\Sigma_{\mathbf{d}}}$, so $\mathscr{O}_{\Sigma_{\mathbf{d}}}$ is locally closed. Smoothness follows because in characteristic 0 the stabilizer $G_{\mathbf{d}}$ is smooth (its identity component is smooth and the component group is finite étale), hence $\operatorname{PGL}_{n}/G_{\mathbf{d}}$ is smooth. ∎
2.3.3. Twisting to the Severi–Brauer scheme
Lemma 2.17 (Full relative Hilbert scheme as an associated bundle).
There is a canonical isomorphism of XX–schemes
$\operatorname{Hilb}\bigl(SB(\mathcal{A})/X\bigr)\ \cong\ P\times^{\operatorname{PGL}_{n}}\operatorname{Hilb}(\mathbb{P}^{n-1}),$
where $\operatorname{PGL}_{n}$ acts on $\operatorname{Hilb}(\mathbb{P}^{n-1})$ via its action on $\mathbb{P}^{n-1}$. Under this identification, universal families correspond by descent.
Proof.
Choose an fppf cover $U\to X$ trivialising the $\operatorname{PGL}_{n}$–torsor $P$, so that
$SB(\mathcal{A})|_{U}\cong\mathbb{P}^{n-1}\times U.$
By base change for Hilbert schemes, for each Hilbert polynomial $\mathsf{P}$ one has
$\operatorname{Hilb}^{\mathsf{P}}\!\bigl(SB(\mathcal{A})/X\bigr)\times_{X}U\ \cong\ \operatorname{Hilb}^{\mathsf{P}}\!\bigl((\mathbb{P}^{n-1}\times U)/U\bigr)\ \cong\ \operatorname{Hilb}^{\mathsf{P}}(\mathbb{P}^{n-1})\times U.$
Taking the disjoint union over all $\mathsf{P}$ yields
$\operatorname{Hilb}\bigl(SB(\mathcal{A})/X\bigr)\times_{X}U\ \cong\ \operatorname{Hilb}(\mathbb{P}^{n-1})\times U.$
On overlaps $U\times_{X}U$, two trivializations differ by a transition function with values in $\operatorname{PGL}_{n}$. The induced gluing on the Hilbert factor is precisely the $\operatorname{PGL}_{n}$–action on $\operatorname{Hilb}(\mathbb{P}^{n-1})$. Thus the gluing rule on overlaps is governed precisely by the $\operatorname{PGL}_{n}$–action on $\operatorname{Hilb}(\mathbb{P}^{n-1})$, which is exactly the associated-bundle construction $P\times^{\operatorname{PGL}_{n}}\operatorname{Hilb}(\mathbb{P}^{n-1})$. ∎
Definition 2.18 (The subsystem-structure locus).
Define the subsystem-structure locus to be the locally closed XX–subscheme
$\operatorname{Hilb}^{\Sigma_{\mathbf{d}}}\!\bigl(SB(\mathcal{A})/X\bigr)\;:=\;P\times^{\operatorname{PGL}_{n}}\mathscr{O}_{\Sigma_{\mathbf{d}}}\ \subset\ P\times^{\operatorname{PGL}_{n}}\operatorname{Hilb}(\mathbb{P}^{n-1})\ \cong\ \operatorname{Hilb}\bigl(SB(\mathcal{A})/X\bigr),$
where $\mathscr{O}_{\Sigma_{\mathbf{d}}}\subset\operatorname{Hilb}(\mathbb{P}^{n-1})$ is the orbit given in Lemma 2.16.
Lemma 2.19 (Identification with the quotient $P/G_{\mathbf{d}}$).
There is a canonical isomorphism of XX–schemes
$\operatorname{Hilb}^{\Sigma_{\mathbf{d}}}\!\bigl(SB(\mathcal{A})/X\bigr)\ \cong\ P/G_{\mathbf{d}},$
and under this identification the inclusion $\operatorname{Hilb}^{\Sigma_{\mathbf{d}}}(SB(\mathcal{A})/X)\hookrightarrow\operatorname{Hilb}(SB(\mathcal{A})/X)$ is a locally closed immersion.
Proof.
Here $P/G_{\mathbf{d}}$ denotes the fppf quotient sheaf, which is represented by an XX–scheme since $G_{\mathbf{d}}\subset\operatorname{PGL}_{n}$ is a smooth affine subgroup. So, $P/G_{\mathbf{d}}\cong P\times^{\operatorname{PGL}_{n}}(\operatorname{PGL}_{n}/G_{\mathbf{d}})$.
By Lemma 2.16, $\mathscr{O}_{\Sigma_{\mathbf{d}}}\cong\operatorname{PGL}_{n}/G_{\mathbf{d}}$ as $\operatorname{PGL}_{n}$–schemes. Hence
$P\times^{\operatorname{PGL}_{n}}\mathscr{O}_{\Sigma_{\mathbf{d}}}\ \cong\ P\times^{\operatorname{PGL}_{n}}(\operatorname{PGL}_{n}/G_{\mathbf{d}}).$
There is a standard isomorphism
$P\times^{\operatorname{PGL}_{n}}(\operatorname{PGL}_{n}/G_{\mathbf{d}})\ \xrightarrow{\sim}\ P/G_{\mathbf{d}},\qquad[p,\ gG_{\mathbf{d}}]\longmapsto(p\cdot g)\,G_{\mathbf{d}},$
which is well-defined and functorial. Local closedness follows fppf-locally on $X$: over a trivialising cover $U\to X$, the inclusion becomes
$(\operatorname{PGL}_{n}/G_{\mathbf{d}})\times U\ \hookrightarrow\ \operatorname{Hilb}(\mathbb{P}^{n-1})\times U,$
which is locally closed by Lemma 2.16, and these local immersions glue. ∎
Proposition 2.20 (Geometric structure of the Hilbert–locus).
Assume $\mathrm{char}(k)=0$, so that $G_{\mathbf{d}}$ is smooth. Then the subsystem-structure locus
$\operatorname{Hilb}^{\Sigma_{\mathbf{d}}}\!\bigl(SB(\mathcal{A})/X\bigr)\ \cong\ P/G_{\mathbf{d}}$
is a smooth $\operatorname{PGL}_{n}/G_{\mathbf{d}}$-fibration. In particular, the structure morphism
$P/G_{\mathbf{d}}\ \longrightarrow\ X$
is smooth of relative dimension $\dim(\operatorname{PGL}_{n}/G_{\mathbf{d}})$.
Moreover, since $G_{\mathbf{d}}^{\circ}$ is (up to finite kernel) the image of $\prod_{i}\operatorname{PGL}_{d_{i}}$, one has
$\dim(\operatorname{PGL}_{n}/G_{\mathbf{d}})\;=\;n^{2}-\sum_{i=1}^{s}d_{i}^{2}+(s-1).$
Proof.
Since $\mathcal{A}$ is an Azumaya algebra, there exists an étale cover $U\to X$ such that $\mathcal{A}|_{U}\simeq M_{n}(\mathcal{O}_{U})$. After such a base change, the associated $\operatorname{PGL}_{n}$–torsor $P\to X$ becomes trivial and one has $P/G_{\mathbf{d}}|_{U}\simeq(\operatorname{PGL}_{n}/G_{\mathbf{d}})\times U$.
In characteristic 0 the stabilizer $G_{\mathbf{d}}$ is smooth, hence the homogeneous space $\operatorname{PGL}_{n}/G_{\mathbf{d}}$ is smooth (cf. Lemma 2.16). Therefore the projection
$(\operatorname{PGL}_{n}/G_{\mathbf{d}})\times U\longrightarrow U$
is smooth of relative dimension $\dim(\operatorname{PGL}_{n}/G_{\mathbf{d}})$, and smoothness (and relative dimension) descends along étale covers. This proves that $P/G_{\mathbf{d}}\to X$ is smooth with the stated relative dimension.
Finally, since $G_{\mathbf{d}}^{\circ}$ is (up to finite kernel) the image of $\prod_{i}\operatorname{PGL}_{d_{i}}$, we have
$\dim(\operatorname{PGL}_{n}/G_{\mathbf{d}})=\dim(\operatorname{PGL}_{n})-\dim(G_{\mathbf{d}})=(n^{2}-1)-\sum_{i=1}^{s}(d_{i}^{2}-1)=n^{2}-\sum_{i=1}^{s}d_{i}^{2}+(s-1),$
and the finite component group does not affect the dimension. ∎
2.3.4. Equivalence of moduli functors and representability
Theorem 2.21 (Subsystem structures as a Hilbert-scheme locus).
Let $\mathcal{A}$ be an Azumaya algebra of degree $n$ on $X$ with associated $\operatorname{PGL}_{n}$–torsor $P\to X$, and fix $\mathbf{d}$ with stabilizer $G_{\mathbf{d}}\subset\operatorname{PGL}_{n}$.
  1. (i)
    The functor $\mathfrak{Red}_{\mathbf{d}}(\mathcal{A})$ of $G_{\mathbf{d}}$–reductions of $P$ is represented by the quotient XX–scheme $P/G_{\mathbf{d}}$.
  2. (ii)
    The subfunctor $\mathfrak{Hilb}^{\Sigma_{\mathbf{d}}}_{\mathcal{A}}\subset\operatorname{Hilb}(SB(\mathcal{A})/X)$ of relative closed subschemes that are fppf-locally isomorphic to $\Sigma_{\mathbf{d}}\times X$ is represented by the same scheme.
  3. (iii)
    Under these identifications, a global $\mathbf{d}$–subsystem structure on $\mathcal{A}$ is equivalently a section
    $X\ \longrightarrow\ P/G_{\mathbf{d}}\ \cong\ \operatorname{Hilb}^{\Sigma_{\mathbf{d}}}\!\bigl(SB(\mathcal{A})/X\bigr).$
Proof.
We construct mutually inverse natural transformations directly, by gluing in families.
(Step 1) From reductions to Segre families. Let $f:T\to X$ and let $(P_{\mathbf{d},T},\alpha_{T})\in\mathfrak{Red}_{\mathbf{d}}(\mathcal{A})(T)$. Form the associated TT–scheme
$\Sigma_{\mathbf{d}}(\mathcal{A}_{T})\ :=\ P_{\mathbf{d},T}\times^{G_{\mathbf{d}}}\Sigma_{\mathbf{d}}.$
Using $\alpha_{T}$, the induced isomorphism of associated $\mathbb{P}^{n-1}$–bundles
$P_{\mathbf{d},T}\times^{G_{\mathbf{d}}}\mathbb{P}^{n-1}\ \xrightarrow{\sim}\ P_{T}\times^{\operatorname{PGL}_{n}}\mathbb{P}^{n-1}\;=\;SB(\mathcal{A}_{T})$
identifies the descended closed immersion $\Sigma_{\mathbf{d}}(\mathcal{A}_{T})\hookrightarrow P_{\mathbf{d},T}\times^{G_{\mathbf{d}}}\mathbb{P}^{n-1}$ (with fiberwise the classical Segre embedding) with a closed immersion
$\Sigma_{\mathbf{d}}(\mathcal{A}_{T})\hookrightarrow SB(\mathcal{A}_{T})=SB(\mathcal{A})\times_{X}T.$
Let $\Sigma_{T}$ denote its image. Then $\Sigma_{T}$ is flat over $T$ and fppf-locally on $T$ it identifies with $\Sigma_{\mathbf{d}}\times T$, hence $\Sigma_{T}\in\mathfrak{Hilb}^{\Sigma_{\mathbf{d}}}_{\mathcal{A}}(T)$. This yields a natural transformation
$\Theta:\ \mathfrak{Red}_{\mathbf{d}}(\mathcal{A})\to\mathfrak{Hilb}^{\Sigma_{\mathbf{d}}}_{\mathcal{A}}.$
(Step 2) From Segre families to reductions. Conversely, let $\Sigma_{T}\in\mathfrak{Hilb}^{\Sigma_{\mathbf{d}}}_{\mathcal{A}}(T)$. Choose an fppf cover $\{T_{i}\to T\}$ satisfying Definition 2.15. Thus $SB(\mathcal{A})\times_{X}T_{i}\simeq\mathbb{P}^{n-1}\times T_{i}$ and $\Sigma_{T}|_{T_{i}}\simeq\Sigma_{\mathbf{d}}\times T_{i}$. On overlaps $T_{ij}$, the transition functions for the $\mathbb{P}^{n-1}$–bundle $SB(\mathcal{A})\times_{X}T\to T$ are maps $g_{ij}:T_{ij}\to\operatorname{PGL}_{n}$. Since the Segre subschemes match in both trivializations, each $g_{ij}$ preserves $\Sigma_{\mathbf{d}}$, hence takes values in $G_{\mathbf{d}}$ by definition of the stabilizer. Therefore $\{g_{ij}\}$ is a $G_{\mathbf{d}}$–valued cocycle, defining a principal $G_{\mathbf{d}}$–torsor $P_{\mathbf{d},T}\to T$ and an induced identification $P_{\mathbf{d},T}\times^{G_{\mathbf{d}}}\operatorname{PGL}_{n}\simeq P_{T}$. This yields a natural transformation
$\Psi:\ \mathfrak{Hilb}^{\Sigma_{\mathbf{d}}}_{\mathcal{A}}\to\mathfrak{Red}_{\mathbf{d}}(\mathcal{A}).$
(Step 3) Inverses. The constructions are inverse because they are governed by the same $G_{\mathbf{d}}$–valued descent datum. If one starts with a reduction, the descended Segre family has transition functions in $G_{\mathbf{d}}$. Conversely, if one starts with a Segre family, the transition functions of $SB(\mathcal{A}_{T})$ are forced to lie in $G_{\mathbf{d}}$, and this recovers the reduction.
(Step 4) Representability. Let $f:T\to X$ be an XX–scheme and let $\Sigma_{T}\subset SB(\mathcal{A})\times_{X}T$ be the family corresponding to a morphism $\xi:T\to\operatorname{Hilb}\bigl(SB(\mathcal{A})/X\bigr)$.
By Lemma 2.17 one has a natural identification
$\operatorname{Hilb}\bigl(SB(\mathcal{A})/X\bigr)\;\simeq\;P\times^{\operatorname{PGL}_{n}}\operatorname{Hilb}(\mathbb{P}^{n-1}),$
and by Definition 2.18 the subsystem-structure locus (Hilbert–locus) is
$\operatorname{Hilb}^{\Sigma_{\mathbf{d}}}\bigl(SB(\mathcal{A})/X\bigr)\;=\;P\times^{\operatorname{PGL}_{n}}\mathscr{O}_{\Sigma_{\mathbf{d}}}\;\subset\;P\times^{\operatorname{PGL}_{n}}\operatorname{Hilb}(\mathbb{P}^{n-1}).$
Since $\mathscr{O}_{\Sigma_{\mathbf{d}}}\hookrightarrow\operatorname{Hilb}(\mathbb{P}^{n-1})$ is a locally closed immersion, it is a monomorphism. Hence the condition that a morphism factors through $P\times^{\operatorname{PGL}_{n}}\mathscr{O}_{\Sigma_{\mathbf{d}}}$ may be checked fppf–locally on $T$.
Choose an fppf covering $\{T_{i}\to T\}$ trivializing the pullback torsor $P_{T}\to T$, so that $SB(\mathcal{A})\times_{X}T_{i}\simeq\mathbb{P}^{n-1}\times T_{i}$. Over each $T_{i}$ the family $\Sigma_{T}|_{T_{i}}$ corresponds to a classifying morphism $\xi_{i}:T_{i}\to\operatorname{Hilb}(\mathbb{P}^{n-1})$. By construction, $\xi$ factors through $P\times^{\operatorname{PGL}_{n}}\mathscr{O}_{\Sigma_{\mathbf{d}}}$ if and only if each $\xi_{i}$ factors through $\mathscr{O}_{\Sigma_{\mathbf{d}}}\subset\operatorname{Hilb}(\mathbb{P}^{n-1})$.
On the other hand, by Definition 2.15, $\Sigma_{T}\in\mathfrak{Hilb}^{\Sigma_{\mathbf{d}}}_{\mathcal{A}}(T)$ means precisely that after an fppf refinement of $\{T_{i}\to T\}$ one may choose trivializations $SB(\mathcal{A})\times_{X}T_{ij}\simeq\mathbb{P}^{n-1}\times T_{ij}$ under which $\Sigma_{T}|_{T_{ij}}$ identifies with $\Sigma_{\mathbf{d}}\times T_{ij}$. After such a refinement the corresponding classifying maps $T_{ij}\to\operatorname{Hilb}(\mathbb{P}^{n-1})$ land in the $\operatorname{PGL}_{n}$–orbit $\mathscr{O}_{\Sigma_{\mathbf{d}}}$. Thus
$\Sigma_{T}\in\mathfrak{Hilb}^{\Sigma_{\mathbf{d}}}_{\mathcal{A}}(T)\quad\Longleftrightarrow\quad\xi\text{ factors through }\operatorname{Hilb}^{\Sigma_{\mathbf{d}}}\!\bigl(SB(\mathcal{A})/X\bigr),$
so $\operatorname{Hilb}^{\Sigma_{\mathbf{d}}}\!\bigl(SB(\mathcal{A})/X\bigr)$ represents the functor $\mathfrak{Hilb}^{\Sigma_{\mathbf{d}}}_{\mathcal{A}}$.
Finally, Lemma 2.19 identifies $\operatorname{Hilb}^{\Sigma_{\mathbf{d}}}\bigl(SB(\mathcal{A})/X\bigr)\cong P/G_{\mathbf{d}}$. Combining this with Steps 2–3, we conclude that $P/G_{\mathbf{d}}$ represents both $\mathfrak{Hilb}^{\Sigma_{\mathbf{d}}}_{\mathcal{A}}$ and $\mathfrak{Red}_{\mathbf{d}}(\mathcal{A})$. Statement (iii) is obtained by evaluating the representing property at $T=X$.
Remark 2.22.
The identification in Theorem 2.21 (iii)
$\operatorname{Hilb}^{\Sigma_{\mathbf{d}}}\bigl(SB(\mathcal{A})/X\bigr)\ \cong\ P/G_{\mathbf{d}}$
realizes the moduli of subsystem structures as a canonically defined locally closed subscheme of the full relative Hilbert scheme. In particular, by Proposition 2.20 it is a smooth XX–scheme (an étale-locally trivial $\operatorname{PGL}_{n}/G_{\mathbf{d}}$–bundle).
Let $\operatorname{Hilb}_{[\Sigma_{\mathbf{d}}]}(SB(\mathcal{A})/X)\subset\operatorname{Hilb}(SB(\mathcal{A})/X)$ denote the open-and-closed component containing the Segre point (the component corresponding to the Hilbert polynomial of $\Sigma_{\mathbf{d}}$). This component is projective over $X$. Hence the closure
$\overline{\operatorname{Hilb}^{\Sigma_{\mathbf{d}}}}\ \subset\ \operatorname{Hilb}_{[\Sigma_{\mathbf{d}}]}\bigl(SB(\mathcal{A})/X\bigr)$
provides a canonical projective compactification of $P/G_{\mathbf{d}}$ over $X$. Its boundary $\partial:=\overline{\operatorname{Hilb}^{\Sigma_{\mathbf{d}}}}\setminus\operatorname{Hilb}^{\Sigma_{\mathbf{d}}}$ parametrizes flat degenerations of product-state loci inside the twisted family $SB(\mathcal{A})\to X$: fiberwise, points of $\partial$ correspond to limits of Segre fibers in $\mathbb{P}^{n-1}$.
In the case $\mathbf{d}=(2,2)$, this boundary becomes explicit. The relevant Hilbert component is $\mathbb{P}^{9}$ (Proposition 2.23), and the boundary is the discriminant divisor of singular quadrics. Twisting yields the corresponding relative discriminant divisor (Corollary 2.24).
2.4. The case $\mathbf{d}=(2,2)$
Let $n=4$ and $\mathbf{d}=(2,2)$. Then $\Sigma_{2,2}\subset\mathbb{P}^{3}$ is a smooth quadric surface. Let $\mathscr{H}:=\operatorname{Hilb}_{[\Sigma_{2,2}]}(\mathbb{P}^{3})$ be the Hilbert scheme component containing $\Sigma_{2,2}$.
Proposition 2.23.
There is a natural isomorphism
$\mathscr{H}\;\simeq\;\mathbb{P}\big(\mathrm{Sym}^{2}(k^{4})^{\vee}\big)\;\simeq\;\mathbb{P}^{9},$
parameterising quadric hypersurfaces in $\mathbb{P}^{3}$. Under this identification, the $\operatorname{PGL}_{4}$-orbit of $\Sigma_{2,2}]$ equals the open subset of smooth quadrics, and its complement is the discriminant hypersurface of singular quadrics. Moreover, the stabilizer of $\Sigma_{2,2}]$ in $\operatorname{PGL}_{4}$ is $G_{(2,2)}$ (so the open orbit is $\operatorname{PGL}_{4}/G_{(2,2)}$).
Proof.
A quadric in $\mathbb{P}^{3}$ is defined by a homogeneous polynomial of degree $2$, hence is a point of $\mathbb{P}(\mathrm{Sym}^{2}(k^{4})^{\vee})=\mathbb{P}^{9}$. The universal family of quadrics is flat, giving a morphism $\mathbb{P}^{9}\to\operatorname{Hilb}(\mathbb{P}^{3})$ whose image is the component $\mathscr{H}$, and this morphism is an isomorphism. Smoothness is detected by the nonvanishing of the determinant of the associated symmetric matrix, so the smooth locus is the complement of the discriminant hypersurface. Transitivity of $\operatorname{PGL}_{4}$ on smooth quadrics identifies this open locus with $\operatorname{PGL}_{4}/G_{(2,2)}$. ∎
Corollary 2.24.
Let $P\to X$ be a $\operatorname{PGL}_{4}$-torsor, and let $SB(\mathcal{A})\to X$ be the associated Severi–Brauer fibration. Then the Hilbert component $\operatorname{Hilb}_{[\Sigma_{2,2}]}(SB(\mathcal{A})/X)$ is the $\mathbb{P}^{9}$-bundle
$P\times^{\operatorname{PGL}_{4}}\mathbb{P}^{9}\;\to\;X,$
and the subsystem-structure space $P/G_{(2,2)}$ identifies with the open complement of the relative discriminant divisor. In particular, the closure of $P/G_{(2,2)}$ inside this Hilbert component is projective over $X$ and its boundary is a divisor (the relative discriminant).
Proof.
Apply the functorial identification of the relative Hilbert scheme with the associated bundle $P\times^{\operatorname{PGL}_{4}}\operatorname{Hilb}(\mathbb{P}^{3})$, and use Proposition 2.23 fiberwise. ∎
3. Entanglement filtration
From here on we assume $\mathbf{d}=(d_{A},d_{B})$ with $n=d_{A}d_{B}$. Write $\Sigma_{A,B}:=\mathbb{P}^{d_{A}-1}\times\mathbb{P}^{d_{B}-1}$ for the product-state locus. In the bipartite case we also write $\Sigma_{A,B}=\Sigma_{(d_{A},d_{B})}$.
Remark 3.1 (Global nature of determinantal geometry in the twisted setting).
Even though the determinantal loci $R_{\leq r}\subset\mathbb{P}(V_{A}\otimes V_{B})$ are classical, they do not make sense on a general Severi–Brauer scheme without extra data. A twisted $\mathbb{P}^{n-1}$–bundle carries no canonical global tensor-product identification. A $\mathbf{d}$–subsystem structure is precisely the additional datum required to globalize the entire entanglement filtration, its incidence resolutions, and the resulting intersection-theoretic invariants uniformly over $X$.
3.1. Entanglement filtration and relative geometry
3.1.1. Schmidt-rank loci as $G_{\mathbf{d}}$–invariant subschemes
Let $V_{A}:=k^{d_{A}}$ and $V_{B}:=k^{d_{B}}$. For $1\leq r\leq\min(d_{A},d_{B})$, let $R_{\leq r}\subset\mathbb{P}(V_{A}\otimes V_{B})$ be the classical determinantal variety of tensors of (matrix) rank $\leq r$, defined by the $(r+1)\times(r+1)$ minors of a flattening matrix. It is a closed, irreducible, $\operatorname{GL}(V_{A})\times\operatorname{GL}(V_{B})$–invariant subscheme.
Definition 3.2 (Loci and strata).
Let $(\mathcal{A},P_{\mathbf{d}})$ be an Azumaya algebra with a chosen subsystem structure. Define the closed subscheme
$\Sigma_{\leq r}(\mathcal{A},\mathbf{d})\;:=\;P_{\mathbf{d}}\times^{G_{\mathbf{d}}}R_{\leq r}\;\subset\;P_{\mathbf{d}}\times^{G_{\mathbf{d}}}\mathbb{P}(V_{A}\otimes V_{B})\;\cong\;SB(\mathcal{A}).$
We call $\Sigma_{\leq r}(\mathcal{A},\mathbf{d})$ the Schmidt-rank $\leq r$ locus, and $\Sigma_{=r}(\mathcal{A},\mathbf{d}):=\Sigma_{\leq r}(\mathcal{A},\mathbf{d})\setminus\Sigma_{\leq r-1}(\mathcal{A},\mathbf{d})$ the Schmidt-rank $r$ stratum when $r>1$. In particular, when $r=1$, they are called product-state locus and product-state stratum, respectively. We then have the natural entanglement filtration:
$\Sigma_{\leq 1}(\mathcal{A},\mathbf{d})\subset\Sigma_{\leq 2}(\mathcal{A},\mathbf{d})\subset\cdots\subset SB(\mathcal{A}).$
By definition, for all $r\geq 2$, every element of $\Sigma_{=r}$ is entangled, while every element of $\Sigma_{=1}$ is a product state. We define $\Sigma_{=0}=\emptyset$.
Proposition 3.3 (Entangled determinantal loci: flatness and base change).
Each $\Sigma_{\leq r}(\mathcal{A},\mathbf{d})\subset SB(\mathcal{A})$ is a well-defined closed subscheme, flat over $X$. Moreover, for any morphism $f:Y\to X$,
$\Sigma_{\leq r}(f^{*}\mathcal{A},\mathbf{d})\;\cong\;f^{*}\Sigma_{\leq r}(\mathcal{A},\mathbf{d})$
as closed subschemes of $SB(f^{*}\mathcal{A})\cong f^{*}SB(\mathcal{A})$.
Proof.
The determinantal variety $R_{\leq r}\subset\mathbb{P}(V_{A}\otimes V_{B})$ is invariant under the stabilizer $G_{\mathbf{d}}\subset\operatorname{PGL}_{n}$. Hence, in a $G_{\mathbf{d}}$–adapted local trivialisation of $SB(\mathcal{A})$, the closed subscheme $R_{\leq r}$ is carried consistently across overlaps by the transition functions. This produces a globally defined closed subscheme
$\Sigma_{\leq r}(\mathcal{A},\mathbf{d})=P_{\mathbf{d}}\times^{G_{\mathbf{d}}}R_{\leq r}\ \subset\ SB(\mathcal{A}).$
Flatness follows because fppf-locally on $X$ the pair $\bigl(SB(\mathcal{A}),\Sigma_{\leq r}(\mathcal{A},\mathbf{d})\bigr)$ is isomorphic to $\bigl(\mathbb{P}(V_{A}\otimes V_{B}),R_{\leq r}\bigr)\times U$, and $R_{\leq r}\times U\to U$ is flat. Compatibility with base change is formal for associated bundles. ∎
Remark 3.4.
The filtration
$\Sigma_{A,B}(\mathcal{A})=\Sigma_{\leq 1}(\mathcal{A},\mathbf{d})\subset\Sigma_{\leq 2}(\mathcal{A},\mathbf{d})\subset\cdots\subset\Sigma_{\leq\min(d_{A},d_{B})}(\mathcal{A},\mathbf{d})=SB(\mathcal{A})$
is the algebro-geometric counterpart of the entanglement hierarchy by Schmidt rank. It exists only when a subsystem structure exists.
3.1.2. Relative singularities, normality, and canonical resolutions
We continue in the bipartite case $\mathbf{d}=(d_{A},d_{B})$ with $n=d_{A}d_{B}$. Let $V_{A}:=k^{d_{A}}$ and $V_{B}:=k^{d_{B}}$, so that $SB(\mathcal{A})_{x}\cong\mathbb{P}(V_{A}\otimes V_{B})$ in a $G_{\mathbf{d}}$–adapted local trivialisation.
Recall that $R_{\leq r}\subset\mathbb{P}(V_{A}\otimes V_{B})$ denotes the classical determinantal variety of tensors (of matrix rank) $\leq r$, and that
[…]
Let
$\operatorname{Mat}_{d_{A}\times d_{B}}:=\operatorname{Hom}(V_{B}^{\vee},V_{A})$
be the affine space of $d_{A}\times d_{B}$ matrices (rows indexed by $V_{A}$, columns indexed by $V_{B}$). Via either natural identification $V_{A}\otimes V_{B}\cong\operatorname{Hom}(V_{A}^{\vee},V_{B})\cong\operatorname{Hom}(V_{B}^{\vee},V_{A})$, matrix rank coincides with tensor (Schmidt) rank in the bipartite model.
Let $Z_{\leq r}\subset\operatorname{Mat}_{d_{A}\times d_{B}}$ be the affine determinantal locus defined by the $(r+1)\times(r+1)$ minors.
Proposition 3.5 (Singular locus of determinantal varieties).
Let $1\leq r\leq\min(d_{A},d_{B})$.
  1. (i)
    If $1\leq r<\min(d_{A},d_{B})$, then the singular locus is
    $\operatorname{Sing}(Z_{\leq r})=Z_{\leq r-1}\qquad\text{and}\qquad\operatorname{Sing}(R_{\leq r})=R_{\leq r-1}.$
  2. (ii)
    If $r=\min(d_{A},d_{B})$, then $R_{\leq r}=\mathbb{P}(V_{A}\otimes V_{B})$ is smooth, hence
    $\operatorname{Sing}(R_{\leq r})=\varnothing.$
Proof.
This is classical determinantal geometry: for the affine determinantal variety $Z_{\leq r}$ of $d_{A}\times d_{B}$ matrices, the singular locus is $Z_{\leq r-1}$ for $r<\min(d_{A},d_{B})$, and $Z_{\leq r}$ is smooth for $r=\min(d_{A},d_{B})$. Projectivizing gives $\operatorname{Sing}(R_{\leq r})=R_{\leq r-1}$ in the projective space. See e.g. standard references on determinantal varieties (e.g. Bruns–Vetter [bruns2006determinantal] or Weyman [Weyman_2003]) for detailed proofs. ∎
There are several standard desingularizations of $R_{\leq r}$. For our relative purposes (and to allow possible discrete symmetries when $d_{A}=d_{B}$), it is convenient to use a symmetric incidence resolution involving both Grassmannians.
Let $\operatorname{Gr}_{A}(r):=\operatorname{Gr}(r,V_{A})$ and $\operatorname{Gr}_{B}(r):=\operatorname{Gr}(r,V_{B})$, with tautological subbundles $\mathcal{U}_{A}\subset V_{A}\otimes\mathcal{O}_{\operatorname{Gr}_{A}(r)}$ and $\mathcal{U}_{B}\subset V_{B}\otimes\mathcal{O}_{\operatorname{Gr}_{B}(r)}$.
Definition 3.6 (Symmetric incidence resolution of $R_{\leq r}$).
Define the smooth projective variety
$\widetilde{R}_{\leq r}\;:=\;\mathbb{P}(\mathcal{U}_{A}\boxtimes\mathcal{U}_{B})\ \longrightarrow\ \operatorname{Gr}_{A}(r)\times\operatorname{Gr}_{B}(r),$
and let
$\rho_{r}:\widetilde{R}_{\leq r}\longrightarrow\mathbb{P}(V_{A}\otimes V_{B})$
be the morphism induced by the natural inclusion $\mathcal{U}_{A}\boxtimes\mathcal{U}_{B}\hookrightarrow(V_{A}\otimes V_{B})\otimes\mathcal{O}$. Its image lies in $R_{\leq r}$, since any tensor in $U_{A}\otimes U_{B}$ has rank $\leq r$.
Proposition 3.7 (Properties of the incidence resolution).
For $1\leq r\leq\min(d_{A},d_{B})$, the morphism
$\rho_{r}:\widetilde{R}_{\leq r}\longrightarrow R_{\leq r}$
is projective and birational. Moreover,
  1. (i)
    $\widetilde{R}_{\leq r}$ is smooth.
  2. (ii)
    $\rho_{r}$ restricts to an isomorphism over the open locus
    $R_{=r}\;:=\;R_{\leq r}\setminus R_{\leq r-1}$
    of tensors of rank exactly $r$. (In particular, if $r<\min(d_{A},d_{B})$ then $R_{=r}$ is the smooth locus of $R_{\leq r}$ by Proposition 3.5.)
  3. (iii)
    Consequently, $\rho_{r}$ is a resolution of singularities of $R_{\leq r}$ (and for $r=\min(d_{A},d_{B})$, a projective birational morphism from a smooth variety onto the already smooth $R_{\leq r}=\mathbb{P}(V_{A}\otimes V_{B})$).
The construction is functorial for the natural $\operatorname{GL}(V_{A})\times\operatorname{GL}(V_{B})$–action. Moreover, when $d_{A}=d_{B}$ it is compatible with the factor-swap involution (it exchanges the two Grassmannian factors $\operatorname{Gr}_{A}(r)$ and $\operatorname{Gr}_{B}(r)$). Hence the construction is equivariant for the full stabilizer $G_{\mathbf{d}}\subset\operatorname{PGL}(V_{A}\otimes V_{B})$.
Proof.
Smoothness of $\widetilde{R}_{\leq r}=\mathbb{P}(\mathcal{U}_{A}\boxtimes\mathcal{U}_{B})$ follows because it is a projective bundle over the smooth base $\operatorname{Gr}_{A}(r)\times\operatorname{Gr}_{B}(r)$. The map $\rho_{r}$ is projective by construction and is birational since it restricts to an isomorphism over the open rank-$r$ locus $R_{=r}$, where the image subspaces $\mathrm{im}(\psi_{A})$ and $\mathrm{im}(\psi_{B})$ vary regularly. Equivariance under $\operatorname{GL}(V_{A})\times\operatorname{GL}(V_{B})$ (and the factor-swap involution when $d_{A}=d_{B}$) is functorial. For details, see standard treatments of incidence resolutions of determinantal varieties. ∎
Remark 3.8 (One-sided resolutions and small resolutions).
There are also standard one-sided incidence resolutions using only $\operatorname{Gr}_{B}(r)$ (image) or only $\operatorname{Gr}_{A}(r)$ (kernel/transpose). The symmetric resolution above dominates both by forgetting one Grassmannian factor. In certain special cases one-sided resolutions can be small. We do not use smallness here.
3.1.3. Relative entangling: singular locus and resolution for $\Sigma_{\leq r}(\mathcal{A},\mathbf{d})$
Definition 3.9 (Relative incidence resolution of Schmidt-rank loci).
Let $(\mathcal{A},P_{\mathbf{d}})$ be an Azumaya algebra with a chosen subsystem structure. Define the XX–scheme
$\widetilde{\Sigma}_{\leq r}(\mathcal{A},\mathbf{d})\;:=\;P_{\mathbf{d}}\times^{G_{\mathbf{d}}}\widetilde{R}_{\leq r},$
and let
$\rho_{r}:\widetilde{\Sigma}_{\leq r}(\mathcal{A},\mathbf{d})\ \longrightarrow\ \Sigma_{\leq r}(\mathcal{A},\mathbf{d})$
be the induced morphism from the $G_{\mathbf{d}}$–equivariant map $\widetilde{R}_{\leq r}\to R_{\leq r}$.
Theorem 3.10 (Relative singular locus and resolution).
Let $X$ be smooth over $k$ (char 0) and let $(\mathcal{A},P_{\mathbf{d}})$ be an Azumaya algebra with a chosen subsystem structure. Fix $1\leq r\leq\min(d_{A},d_{B})$. Then:
  1. (i)
    Relative singular locus: If $1\leq r<\min(d_{A},d_{B})$, then scheme-theoretically
    $\operatorname{Sing}\bigl(\Sigma_{\leq r}(\mathcal{A},\mathbf{d})\bigr)\;=\;\Sigma_{\leq r-1}(\mathcal{A},\mathbf{d}).$
    In particular, let $\pi_{r}:=\pi|_{{\Sigma_{\leq r}(\mathcal{A},\mathbf{d})}}$. The morphism $\pi_{r}:\Sigma_{\leq r}(\mathcal{A},\mathbf{d})\to X$ is smooth exactly over the open stratum $\Sigma_{=r}(\mathcal{A},\mathbf{d}):=\Sigma_{\leq r}(\mathcal{A},\mathbf{d})\setminus\Sigma_{\leq r-1}(\mathcal{A},\mathbf{d})$.
    If $r=\min(d_{A},d_{B})$, then $\Sigma_{\leq r}(\mathcal{A},\mathbf{d})=SB(\mathcal{A})$ is smooth over $X$, hence
    $\operatorname{Sing}\bigl(\Sigma_{\leq r}(\mathcal{A},\mathbf{d})\bigr)=\varnothing.$
    (Here, one may adopt the convention $\Sigma_{\leq 0}(\mathcal{A},\mathbf{d})=\varnothing$ so that for $r=1$ the formula reads $\operatorname{Sing}(\Sigma_{\leq 1})=\Sigma_{\leq 0}=\varnothing$.)
  2. (ii)
    Resolution: The total space $\widetilde{\Sigma}_{\leq r}(\mathcal{A},\mathbf{d})$ is smooth over $X$, and
    $\rho_{r}:\widetilde{\Sigma}_{\leq r}(\mathcal{A},\mathbf{d})\longrightarrow\Sigma_{\leq r}(\mathcal{A},\mathbf{d})$
    is a projective birational morphism which is an isomorphism over $\Sigma_{=r}(\mathcal{A},\mathbf{d})$. In particular, $\rho_{r}$ gives a resolution of singularities of $\Sigma_{\leq r}(\mathcal{A},\mathbf{d})$.
  3. (iii)
    Base change compatibility: For any morphism $f:Y\to X$,
    $f^{*}\widetilde{\Sigma}_{\leq r}(\mathcal{A},\mathbf{d})\ \cong\ \widetilde{\Sigma}_{\leq r}(f^{*}\mathcal{A},\mathbf{d}),\qquad f^{*}\Sigma_{\leq r}(\mathcal{A},\mathbf{d})\ \cong\ \Sigma_{\leq r}(f^{*}\mathcal{A},\mathbf{d}),$
    and these identifications intertwine the resolution maps.
Proof.
Choose an fppf cover (in fact étale since $G_{\mathbf{d}}$ is smooth) $U\to X$ trivialising $P_{\mathbf{d}}$. After choosing a trivialization of $P_{\mathbf{d}}|_{U}$, we obtain isomorphisms
$SB(\mathcal{A})|_{U}\ \cong\ \mathbb{P}(V_{A}\otimes V_{B})\times U,\qquad\Sigma_{\leq r}(\mathcal{A},\mathbf{d})|_{U}\ \cong\ R_{\leq r}\times U,$
and similarly
$\widetilde{\Sigma}_{\leq r}(\mathcal{A},\mathbf{d})|_{U}\ \cong\ \widetilde{R}_{\leq r}\times U,\qquad\rho_{r}|_{U}\ =\ (\rho_{r}:\widetilde{R}_{\leq r}\to R_{\leq r})\times\mathrm{id}_{U}.$
(i) Singular locus. Assume first that $1\leq r<\min(d_{A},d_{B})$. By Proposition 3.5, $\operatorname{Sing}(R_{\leq r})=R_{\leq r-1}$. Since $U$ is smooth, taking the product with $U$ preserves singular loci:
$\operatorname{Sing}(R_{\leq r}\times U)\;=\;\operatorname{Sing}(R_{\leq r})\times U\;=\;R_{\leq r-1}\times U.$
Indeed, the singular locus is defined by scheme-theoretically by the appropriate Fitting ideal of $\Omega^{1}$, and for $U$ smooth one has $\Omega^{1}_{R_{\leq r}\times U/k}\cong\mathrm{pr}_{1}^{*}\Omega^{1}_{R_{\leq r}/k}\oplus\mathrm{pr}_{2}^{*}\Omega^{1}_{U/k}$, so the Fitting ideal pulls back from $R_{\leq r}$.
Transporting back via the identifications gives
$\operatorname{Sing}\bigl(\Sigma_{\leq r}(\mathcal{A},\mathbf{d})|_{U}\bigr)\;=\;\Sigma_{\leq r-1}(\mathcal{A},\mathbf{d})|_{U}.$
Because $\operatorname{Sing}(-)$ is defined scheme-theoretically by the relevant Fitting ideal of $\Omega^{1}$ and Fitting ideals commute with étale base change, the equality descends, hence $\operatorname{Sing}(\Sigma_{\leq r}(\mathcal{A},\mathbf{d}))=\Sigma_{\leq r-1}(\mathcal{A},\mathbf{d})$ scheme-theoretically on $X$. The equivalent smoothness statement for $\pi_{r}$ follows immediately.
If $r=\min(d_{A},d_{B})$, then $R_{\leq r}=\mathbb{P}(V_{A}\otimes V_{B})$ is smooth, so $\Sigma_{\leq r}(\mathcal{A},\mathbf{d})|_{U}\cong\mathbb{P}(V_{A}\otimes V_{B})\times U$ is smooth. Hence $\Sigma_{\leq r}(\mathcal{A},\mathbf{d})=SB(\mathcal{A})$ is smooth over $X$ and its singular locus is empty.
(ii) Resolution and smoothness over $X$. By Proposition 3.7, $\widetilde{R}_{\leq r}$ is smooth and $\rho_{r}:\widetilde{R}_{\leq r}\to R_{\leq r}$ is a projective birational morphism, an isomorphism over $R_{=r}$. Therefore $\widetilde{R}_{\leq r}\times U$ is smooth over $U$, and the product map $(\rho_{r}\times\mathrm{id}_{U})$ is a resolution of $R_{\leq r}\times U$. By descent (again using that smoothness and properness are fpqc-local on the source/target), $\widetilde{\Sigma}_{\leq r}(\mathcal{A},\mathbf{d})$ is smooth over $X$, and $\rho_{r}$ is projective birational and an isomorphism over $\Sigma_{=r}$.
(iii) Base change. Both $\Sigma_{\leq r}$ and $\widetilde{\Sigma}_{\leq r}$ are defined as associated bundles $P_{\mathbf{d}}\times^{G_{\mathbf{d}}}(\,\cdot\,)$, hence commute with pullback in the base. The compatibility of the maps follows formally from functoriality of associated bundles. ∎
Corollary 3.11.
For $1\leq r<\min(d_{A},d_{B})$, the locus $\Sigma_{\leq r}(\mathcal{A},\mathbf{d})$ is normal and Cohen–Macaulay. Moreover, in characteristic 0 it has rational singularities (hence in particular is Cohen–Macaulay).
Proof.
These properties are étale local on the source. Over an étale cover $U\to X$ trivialising $P_{\mathbf{d}}$ one has $\Sigma_{\leq r}|_{U}\cong R_{\leq r}\times U$. The classical determinantal variety $R_{\leq r}$ is normal and Cohen–Macaulay (and has rational singularities in char 0), and taking the product with a smooth scheme preserves these properties. Therefore $\Sigma_{\leq r}(\mathcal{A},\mathbf{d})$ has the same local singularity profile. ∎
Remark 3.12.
The resolution $\rho_{r}$ is canonical once the subsystem structure is fixed, and it exists uniformly in families. Thus the entanglement filtration supports a range of birational/intersection-theoretic tools (resolutions, exceptional loci, pushforwards, etc.) on the Severi–Brauer background.
3.2. Cycle classes and computability in the split model
The strongest universal output of the construction is the cycle class
$\Sigma_{\leq r}(\mathcal{A},\mathbf{d})]\in CH^{*}(SB(\mathcal{A})).$
This cycle class is not intrinsic to $\mathcal{A}$, but becomes canonical once a subsystem structure $\mathbf{d}$ is chosen and fixed.
Corollary 3.13 (A computable numerical invariant in twisted families).
Assume $X$ is connected and smooth, and let $(\mathcal{A},P_{\mathbf{d}})$ be equipped with a bipartite subsystem structure $\mathbf{d}=(d_{A},d_{B})$. Let $\pi:SB(\mathcal{A})\to X$ and write $H:=c_{1}(\mathcal{O}_{SB(\mathcal{A})}(1))$. For $1\leq r\leq\min(d_{A},d_{B})$ set
$N_{r}:=\dim(R_{\leq r})=r(d_{A}+d_{B}-r)-1.$
Then the class
$\pi_{*}\bigl(H^{N_{r}}\cap[\Sigma_{\leq r}(\mathcal{A},\mathbf{d})]\bigr)\ \in\ CH^{0}(X)$
is the constant integer $\deg(R_{\leq r})$ on each connected component of $X$. So every geometric fiber $(\Sigma_{\leq r})_{x}\subset SB(\mathcal{A})_{x}\simeq\mathbb{P}^{n-1}$ has the same projective degree $\deg(R_{\leq r})$ as the classical determinantal variety.
Proof.
The statement is fiberwise and follows from flatness. Indeed, $\Sigma_{\leq r}(\mathcal{A},\mathbf{d})\to X$ is flat with fibers isomorphic (fppf-locally on $X$) to the fixed projective subscheme $R_{\leq r}\subset\mathbb{P}^{n-1}$ equipped with the hyperplane polarization. Hence the Hilbert polynomial of the fibers of $\Sigma_{\leq r}(\mathcal{A},\mathbf{d})\to X$ with respect to $\mathcal{O}_{SB(\mathcal{A})}(1)|_{\Sigma_{\leq r}}$ is locally constant, and therefore constant on each connected component of $X$.
By definition, the projective degree of a fiber is the leading coefficient of this Hilbert polynomial (up to the standard factorial normalization), so the fiberwise degree is constant on $X$. On a geometric point $x\to X$ where the subsystem structure trivializes, the fiber $(\Sigma_{\leq r})_{x}\subset SB(\mathcal{A})_{x}\simeq\mathbb{P}^{n-1}$ identifies with $R_{\leq r}$, hence has degree $\deg(R_{\leq r})$. Therefore every fiber has degree $\deg(R_{\leq r})$.
Finally, the class $\pi_{*}\!\bigl(H^{N_{r}}\cap[\Sigma_{\leq r}(\mathcal{A},\mathbf{d})]\bigr)\in CH^{0}(X)$ is determined by these fiberwise degrees, and equals $\deg(R_{\leq r})\cdot[X]$ on each connected component. ∎
Example 3.14 (A twisted $(p,p)$ model with an explicit relative degree).
Work with the setup of Proposition 4.16 on $X=\mathbb{G}_{\mathrm{m}}^{2}$, so that the degree-$p^{2}$ Azumaya algebra $A$ admits a global (p,p)(p,p)–subsystem structure. Then the product-state locus $\Sigma_{\leq 1}(A,(p,p))\subset SB(A)$ is a flat family whose geometric fibers are the Segre varieties $\mathbb{P}^{p-1}\times\mathbb{P}^{p-1}\subset\mathbb{P}^{p^{2}-1}$. Hence every fiber has degree
$\deg\bigl((\Sigma_{\leq 1})_{x}\bigr)=\deg\bigl(\mathbb{P}^{p-1}\times\mathbb{P}^{p-1}\bigr)=\binom{2p-2}{p-1}$.
Consequently, with $H=c_{1}(\mathcal{O}_{SB(A)}(1))$ one has in $CH^{0}(X)$
$\pi_{*}\bigl(H^{2p-2}\cap[\Sigma_{\leq 1}(A,(p,p))]\bigr)\;=\;\binom{2p-2}{p-1}\cdot[X].$
For $p=2$ this recovers the familiar degree-$2$ quadric surface in $\mathbb{P}^{3}$.
In the split case one can compute these classes by standard determinantal formulas.
Proposition 3.15 (Split computation via Thom–Porteous).
Assume $\mathcal{A}\cong\underline{\operatorname{End}}(E)$ is split and the subsystem structure comes from a tensor factorization $E\cong E_{A}\otimes E_{B}$ with $\mathrm{rk}(E_{A})=d_{A}$, $\mathrm{rk}(E_{B})=d_{B}$. Then $SB(\mathcal{A})\cong\mathbb{P}(E)$ and $\Sigma_{\leq r}(\mathcal{A},\mathbf{d})$ is the degeneracy locus of the universal map
$\Phi:\ \pi^{*}(E_{A}^{\vee})\otimes\mathcal{O}_{\mathbb{P}(E)}(-1)\longrightarrow\pi^{*}(E_{B}),$
and its class in $CH^{*}(\mathbb{P}(E))$ is given by the Thom–Porteous formula in terms of the Chern classes of the virtual bundle $\pi^{*}(E_{B})-\pi^{*}(E_{A}^{\vee})\otimes\mathcal{O}_{\mathbb{P}(E)}(-1)$.
Proof.
A point $\psi]\in\mathbb{P}(E_{x})=\mathbb{P}(E_{A,x}\otimes E_{B,x})$ corresponds to a tensor $\psi$, hence a linear map $E_{A,x}^{\vee}\to E_{B,x}$ by contraction. The rank of this map is exactly the Schmidt rank of $\psi$. Globalising over $\mathbb{P}(E)$ uses the tautological line subbundle $\mathcal{O}(-1)\subset\pi^{*}E$ and yields $\Phi$. The degeneracy locus $\{\mathrm{rank}(\Phi)\leq r\}$ is precisely $\Sigma_{\leq r}$, and Thom–Porteous gives the class. ∎
Remark 3.16.
Even in the split case, the determinantal strata depend on the chosen tensor factorization $E\cong E_{A}\otimes E_{B}$. There is no canonical way to extract such a factorization from $E$ or from $\underline{\operatorname{End}}(E)$. Thus, the strata and their classes become new geometry created by the subsystem (entanglement) structure.
4. Examples: explicit obstruction computations and non-Brauer invariance
Summary (Outline of the examples).
  1. (i)
    Brauer-theoretic obstruction (torsion). Proposition 4.1 gives the general $\mathrm{lcm}(\mathbf{d})$–torsion constraint, and Theorem 4.7 exhibits a concrete Kummer $p^{2}$–symbol class on $X=\mathbb{G}_{\mathrm{m}}^{2}$ that cannot admit a global $(p,p)$ subsystem structure.
  2. (ii)
    Twisted but reducible case. Proposition 4.16 constructs a nontrivial twisted degree-$p^{2}$ class on the same base that does globalize a $(p,p)$ subsystem structure (hence supports the full entanglement filtration). Example 3.14 records an explicit numerical invariant of its product-state locus.
  3. (iii)
    Torsion is not sufficient. Theorem 4.20 shows that even exponent-$p$ classes can fail to lie in $\operatorname{Br}_{(p,p)}$: the obstruction depends on finer torsor data (index/indecomposability phenomena).
  4. (iv)
    Not a Brauer invariant (even split). Proposition 4.27 and Theorem 4.30 show that $\mathbf{d}$–reducibility depends on the underlying $\operatorname{PGL}_{n}$–torsor, not merely on the Brauer class. Example 4.33 gives an explicit split illustration on $\mathbb{P}^{1}$.
  5. (v)
    Quantum-information witness via monodromy. The analytic monodromy picture in §4.2 makes the failure of reduction visible as an entangling holonomic gate. In particular Theorem 4.10 identifies the $p=2$ monodromy with a $\mathrm{CNOT}$-type operation. Example 4.14 collects these phenomena in a single 88–level test case.
  6. (vi)
    Quantum spin-chain. A simple but new quantum spin chain example yielding entanglement by gluing (locally product) ground states of locally defined Hamiltonian on a torus is addressed in Appendix A.
4.1. Obstructions: general constraints and explicit symbol examples
4.1.1. A general period obstruction: an $\mathrm{lcm}(\mathbf{d})$–torsion constraint
Proposition 4.1 (A general period obstruction for $\mathbf{d}$–decomposable Brauer classes).
Fix a subsystem type $\mathbf{d}=(d_{1},\dots,d_{s})$ and set $n=\prod_{i}d_{i}$ and
$\ell:=\operatorname{lcm}(d_{1},\dots,d_{s}).$
Let $G_{\mathbf{d}}\subset\operatorname{PGL}_{n}$ be the stabilizer (Definition 2.3), possibly disconnected when some $d_{i}$ coincide, and let $\operatorname{Br}_{\mathbf{d}}(X)$ be as in Definition 2.7. Then for every scheme $X$ one has
$\operatorname{Br}_{\mathbf{d}}(X)\ \subset\ \operatorname{Br}(X)[\ell].$
In particular, if $\beta\in\operatorname{Br}_{\mathbf{d}}(X)$, then $\operatorname{per}(\beta)\mid\ell$.
Proof.
Let $G_{\mathbf{d}}\subset\operatorname{PGL}_{n}$ be the stabilizer. Define $\widetilde{G}_{\mathbf{d}}$ to be the subgroup fppf sheaf of $\operatorname{GL}_{n}$ generated by
  1. (a)
    the image of the tensor-product homomorphism $\prod_{i=1}^{s}\operatorname{SL}_{d_{i}}\to\operatorname{SL}_{n}\subset\operatorname{GL}_{n}$, and
  2. (b)
    the permutation-matrix lifts of all permutations of equal Segre factors acting on $k^{d_{1}}\otimes\cdots\otimes k^{d_{s}}$.
By construction, the composite morphism of fppf sheaves
$\widetilde{G}_{\mathbf{d}}\hookrightarrow\operatorname{GL}_{n}\twoheadrightarrow\operatorname{PGL}_{n}$
has image $G_{\mathbf{d}}$. In particular, $\widetilde{G}_{\mathbf{d}}\to G_{\mathbf{d}}$ is surjective as an fppf sheaf.
Claim. The kernel of $\widetilde{G}_{\mathbf{d}}\to G_{\mathbf{d}}$ is the scalar subgroup $\mu_{\ell}\subset\operatorname{GL}_{n}$, where $\ell=\mathrm{lcm}(d_{1},\dots,d_{s})$.
Indeed, the kernel consists of those elements mapping to the identity in $\operatorname{PGL}_{n}$, hence of scalar matrices in $\operatorname{GL}_{n}$. Moreover, if an element of $\widetilde{G}_{\mathbf{d}}$ maps to a scalar matrix in $\operatorname{GL}_{n}$, then its image in the (permutation) component group is necessarily trivial (a nontrivial permutation matrix cannot be scalar), so such a scalar element must lie in the tensor-product part. The only scalars in the image of $\prod_{i}\operatorname{SL}_{d_{i}}\to\operatorname{GL}_{n}$ are tensor products of central elements $\zeta_{i}\cdot I_{d_{i}}$ with $\zeta_{i}\in\mu_{d_{i}}$. Such an element maps to $(\prod_{i}\zeta_{i})\cdot I_{n}$. Since $d_{i}\mid\ell$ for all $i$, these scalars lie in $\mu_{\ell}$, and the subgroup they generate is exactly $\mu_{\ell}$. This proves the claim.
Thus we obtain a central extension of fppf sheaves of groups
$1\longrightarrow\mu_{\ell}\longrightarrow\widetilde{G}_{\mathbf{d}}\longrightarrow G_{\mathbf{d}}\longrightarrow 1,$
and a commutative diagram of central extensions
[…]
Let $P_{\mathbf{d}}]\in H^{1}(X,G_{\mathbf{d}})$ and set $P]:=i_{*}[P_{\mathbf{d}}]\in H^{1}(X,\operatorname{PGL}_{n})$. Since $\mu_{\ell}$ is central, the boundary map for the top extension lands in the abelian group
$\partial_{\ell}:H^{1}(X,G_{\mathbf{d}})\longrightarrow H^{2}(X,\mu_{\ell}).$
By functoriality of boundary maps for morphisms of extensions, we have
$\delta_{n}([P])\;=\;(\mu_{\ell}\hookrightarrow\mathbb{G}_{\mathrm{m}})_{*}\bigl(\partial_{\ell}([P_{\mathbf{d}}])\bigr)\ \in\ H^{2}(X,\mathbb{G}_{\mathrm{m}}),$
where $\delta_{n}:H^{1}(X,\operatorname{PGL}_{n})\to H^{2}(X,\mathbb{G}_{\mathrm{m}})$ is the usual connecting morphism.
Finally, the image of $H^{2}(X,\mu_{\ell})\to H^{2}(X,\mathbb{G}_{\mathrm{m}})$ is $\ell$–torsion (because $\mu_{\ell}\subset\mathbb{G}_{\mathrm{m}}$ is killed by $\ell$), so $\ell\cdot\delta_{n}([P])=0$. Hence $\delta_{n}([P])\in\operatorname{Br}(X)[\ell]$, proving $\operatorname{Br}_{\mathbf{d}}(X)\subset\operatorname{Br}(X)[\ell]$. ∎
4.1.2. Period/index obstruction via a Kummer symbol (generic period $m$)
Fix an integer $m\geq 2$ and assume $\mu_{m}\subset k$ (so $k$ contains a primitive $m$-th root of unity $\zeta_{m}$). (Throughout the paper $\mathrm{char}(k)=0$, so $m$ is automatically invertible in $k$.) Let
$X:=\mathbb{G}_{m,k}^{2}=\operatorname{Spec}k[u^{\pm 1},v^{\pm 1}]$
which is smooth, integral, and regular.
Definition 4.2 (A symbol Azumaya algebra of degree $m$).
Define the sheaf of $\mathcal{O}_{X}$–algebras $\mathcal{A}_{(u,v),m}$ on $X$ by generators and relations: it is locally free of rank $m^{2}$ as an $\mathcal{O}_{X}$–module, generated by $x,y$ with relations
$x^{m}=u,\qquad y^{m}=v,\qquad yx=\zeta_{m}\,xy,$
and with the obvious $\mathcal{O}_{X}$–linearity.
For the relationship between symbol constructions and Brauer classes via the norm-residue homomorphism (and its consequences for central simple/Azumaya algebras), see [1983IzMat..21..307M].
Proposition 4.3 (Azumaya and Brauer class).
$\mathcal{A}_{(u,v),m}$ is an Azumaya algebra of degree $m$ on $X$. Its Brauer class $\beta_{(u,v),m}:=[\mathcal{A}_{(u,v),m}]\in\operatorname{Br}(X)$ is mm–torsion.
Proof.
This is the standard Kummer/cyclic (symbol) construction. Over the Kummer cover adjoining an mm–th root of $u$ (and similarly for $v$), the algebra becomes a matrix algebra. Therefore, it is a crossed product algebra associated to the $\mu_{m}$–torsors defined by $u$ and $v$. Hence it an Azumaya algebra over $X$, and its Brauer class is annihilated by $m$. ∎
Definition 4.4 (Period and index).
Let $X$ be integral with function field $K=k(X)$ and let $\beta\in\operatorname{Br}(X)$.
  • The period $\mathrm{per}(\beta)$ is the order of $\beta$ in $\operatorname{Br}(X)$.
  • The index $\mathrm{ind}(\beta)$ is the index of the class $\beta_{K}\in Br(K)$ in the Brauer group of the field $K$ (the minimal degree of a central simple algebra representing $\beta_{K}$).
One always has $\mathrm{per}(\beta)\mid\mathrm{ind}(\beta)$.
Lemma 4.5 (Generic symbol has period $m$).
Let $K=k(u,v)$ be the rational function field. The symbol class $\beta_{(u,v),m}|_{\operatorname{Spec}K}\in\operatorname{Br}(K)$ has period $m$. Consequently, $\operatorname{per}(\beta_{(u,v),m})=m$ in $\operatorname{Br}(X)$, hence also $\mathrm{ind}(\beta_{(u,v),m})=m$.
Proof.
Over $K$ the class is the usual symbol algebra $(u,v)_{m}$. Consider the discrete valuation $v_{u}$ corresponding to the prime divisor $u=0$ on $\operatorname{Spec}k[u,v]$. For $m$ invertible in $k$ and $\mu_{m}\subset k$, the residue map for mm–torsion Brauer classes sends the symbol $(u,v)_{m}$ to the Kummer class of $v$ in the residue field $k(v)$. Since $v$ is not an $m$-th power in $k(v)$, this residue has exact order $m$. Hence $(u,v)_{m}$ has exact order $m$ in $\operatorname{Br}(K)$.
Since $X$ is regular, the natural map $\operatorname{Br}(X)\hookrightarrow\operatorname{Br}(K)$ is injective, so $\operatorname{per}(\beta_{(u,v),m})=m$ in $\operatorname{Br}(X)$ as well. Finally, for a degree-$m$ central simple algebra one has $\mathrm{ind}\mid m$, while $\operatorname{per}=m$ forces $\mathrm{ind}=m$. ∎
4.1.3. Obstruction for the bipartite type $(p,p)$
We now compare this class to the bipartite subsystem type $\mathbf{d}=(p,p)$ (so $n=p^{2}=m$). From this point on we specialize to $m=p^{2}$ (with $p$ prime) in order to discuss the bipartite subsystem type $(p,p)$.
Corollary 4.6 (A necessary torsion constraint for $\mathbf{d}=(p,p)$).
Let $p$ be a prime and let $X$ be any scheme. If $\beta\in\operatorname{Br}_{(p,p)}(X)$, then $\operatorname{per}(\beta)\mid p$.
Proof.
This is Proposition 4.1 with $\ell=\mathrm{lcm}(p,p)=p$. ∎
Theorem 4.7 (A concrete obstruction class not in $\operatorname{Br}_{(p,p)}(X)$).
With $X=\mathbb{G}_{m}^{2}$ and $\beta=\beta_{(u,v),p^{2}}$ as above, one has
$\beta\notin\operatorname{Br}_{(p,p)}(X).$
Proof.
By Lemma 4.5, $\mathrm{per}(\beta)=p^{2}$. But Corollary 4.6 shows that any class in $\operatorname{Br}_{(p,p)}(X)$ must have period dividing pp. Hence $\beta\notin\operatorname{Br}_{(p,p)}(X)$. ∎
Remark 4.8 (Interpretation of the example).
This is an explicit geometric instance of the intrinsic entanglement obstruction: $SB(\mathcal{A}_{(u,v),p^{2}})\to\mathbb{G}_{m}^{2}$ exists, but no bipartite (p,p)(p,p)–entanglement geometry (Segre subfibration / entanglement filtration) can exist for this Brauer class.
4.2. A concrete quantum information system and an entangling monodromy
The results of §2–§5 are algebro–geometric and do not require analytic topology. In this subsection we specialize to $k=\mathbb{C}$ and interpret local trivializations of the $\operatorname{PGL}_{n}$–torsor on $X^{\mathrm{an}}$ in terms of analytic continuation along loops. The resulting monodromy elements in $\operatorname{PGL}_{n}(\mathbb{C})$ provide an intuitive translation of the obstruction no reduction to $G_{\mathbf{d}}$ into the language of quantum operations.
Fix a prime $p$ and set $m=p^{2}$. Put
$\mathcal{X}=(\mathbb{C}^{\times})^{2},\qquad(u,v)\in\mathcal{X},$
and let
$\zeta:=e^{2\pi i/m}.$
We describe an explicit local model of the symbol Azumaya algebra and show that a gauge monodromy becomes entangling with respect to the bipartite type $(p,p)$.
We first define a local (analytic) trivialization and a parameter-dependent representation. Choose a simply connected open set $U\subset\mathcal{X}$ on which we can pick analytic branches of the $m$-th roots $u^{1/m}$ and $v^{1/m}$. Fix the Hilbert space
$\mathcal{H}:=\mathbb{C}^{m},$
with computational basis $\{|r\rangle\}_{r\in\mathbb{Z}/m\mathbb{Z}}$. Define the Weyl (shift/clock) operators on $\mathcal{H}$ by
(4.1) $\mathbf{X}|r\rangle=|r+1\rangle,\qquad\mathbf{Z}|r\rangle=\zeta^{\,r}|r\rangle.$
Then $\mathbf{Z}\mathbf{X}=\zeta\,\mathbf{X}\mathbf{Z}$. For $(u,v)\in U$, define
$x(u,v):=u^{1/m}\mathbf{X},\qquad y(u,v):=v^{1/m}\mathbf{Z}.$
A direct computation gives
$x(u,v)^{m}=u\,\mathrm{id}_{\mathcal{H}},\qquad y(u,v)^{m}=v\,\mathrm{id}_{\mathcal{H}},\qquad y(u,v)\,x(u,v)=\zeta\,x(u,v)\,y(u,v),$
corresponding to the setup defined by Definition 4.2. Thus, on $U$, the generators $x,y$ of the symbol algebra are realized concretely as the $(u,v)$-dependent operators $x(u,v),y(u,v)\in\mathrm{End}(\mathcal{H})$.
Let $(u_{0},v_{0})\in U$ be a base point and consider the loops
$\gamma_{u}(t)=(u_{0}e^{2\pi it},\,v_{0}),\qquad\gamma_{v}(t)=(u_{0},\,v_{0}e^{2\pi it})\qquad(0\leq t\leq 1).$
Analytic continuation of the chosen branches yields
$u^{1/m}\longmapsto\zeta\,u^{1/m}\quad\text{along }\gamma_{u},\qquad v^{1/m}\longmapsto\zeta\,v^{1/m}\quad\text{along }\gamma_{v}.$
Hence $x(u,v)$ picks up a factor $\zeta$ along $\gamma_{u}$, and $y(u,v)$ picks up a factor $\zeta$ along $\gamma_{v}$. These phase changes are absorbed by conjugation:
$y\,x\,y^{-1}=\zeta\,x,\qquad x^{-1}y\,x=\zeta\,y,$
so the corresponding projective monodromies can be taken as
$g_{u}=[\,y(u_{0},v_{0})\,]\in\operatorname{PGL}(\mathcal{H}),\qquad g_{v}=[\,x(u_{0},v_{0})^{-1}\,]\in\operatorname{PGL}(\mathcal{H}).$
Here $\cdot$] denotes the class in $\operatorname{PGL}(\mathcal{H})=\operatorname{GL}(\mathcal{H})/\mathbb{C}^{\times}$.
Moreover, the commutator of lifts in $\operatorname{GL}(\mathcal{H})$ is a scalar:
$g_{u}g_{v}g_{u}^{-1}g_{v}^{-1}\ \sim\ (\mathbf{Z})(\mathbf{X}^{-1})(\mathbf{Z}^{-1})(\mathbf{X})=\zeta^{-1}\mathrm{id}_{\mathcal{H}}.$
Thus the monodromies commute in $\operatorname{PGL}(\mathcal{H})$ but have a nontrivial scalar discrepancy of order $m$ in $\operatorname{GL}(\mathcal{H})$, which is the local manifestation of the Brauer twisting.
We explicitly construct a local bipartite identification $H\simeq\mathbb{C}^{p}\otimes\mathbb{C}^{p}$. Write each $r\in\mathbb{Z}/m\mathbb{Z}$ uniquely as
$r=a+pb,\qquad a,b\in\{0,1,\dots,p-1\}.$
This gives a local identification
$\mathcal{H}\ \cong\ \mathcal{H}_{A}\otimes\mathcal{H}_{B},\qquad\mathcal{H}_{A}\cong\mathbb{C}^{p},\ \mathcal{H}_{B}\cong\mathbb{C}^{p},$
by sending $|a\rangle_{A}\otimes|b\rangle_{B}\leftrightarrow|a+pb\rangle$. In this tensor-product picture, the projective monodromy around $\gamma_{v}$ acts as $\mathbf{X}^{-1}$ (up to an overall phase), so
$g_{v}\sim\mathbf{X}^{-1}.$
Its action on the tensor basis is
$\mathbf{X}^{-1}|a,b\rangle=\begin{cases}|a-1,\ b\rangle&(a\neq 0),\\[4.0pt] |p-1,\ b-1\rangle&(a=0),\end{cases}$
hence the action on the tensor basis depends on $a$, so it cannot be of the form $U_{A}\otimes U_{B}$.
Now consider the product state
$|\psi\rangle:=(|0\rangle_{A}+|1\rangle_{A})\otimes|0\rangle_{B}.$
Applying $\mathbf{X}^{-1}$ gives
$\mathbf{X}^{-1}|\psi\rangle=|0\rangle_{A}\otimes|0\rangle_{B}+|p-1\rangle_{A}\otimes|p-1\rangle_{B}.$
This state has Schmidt rank $2$ (its coefficient matrix has two independent diagonal entries), hence it is entangled. Therefore the monodromy $g_{v}$ is entangling with respect to the local bipartite type $(p,p)$. Thus $g_{v}$ does not lie in the stabilizer (the local-operation group) $\operatorname{PGL}_{p}\times\operatorname{PGL}_{p}$.
Remark 4.9 (The case $p=2$).
When $p=2$, the above computation gives
$(|0\rangle+|1\rangle)\otimes|0\rangle\ \longmapsto\ |00\rangle+|11\rangle,$
which is (up to normalization) the Bell state. This illustrates concretely that the gauge monodromy generated by the $(u,v)$-twist necessarily contains an entangling component.
Theorem 4.10 (Universality of $p=2$ monodromy gate).
Let $\mathcal{X}:=(\mathbb{C}^{\times})^{2}$ with coordinates $(u,v)$, set $m=4$ and $\zeta=e^{2\pi i/m}=i$, and let $\mathbf{X},\mathbf{Z}\in\operatorname{GL}(\mathcal{H})$ ($\mathcal{H}=\mathbb{C}^{4}$) be the Weyl operators as before. Then:
  1. (i)
    $g_{u}$ is local: in $\operatorname{PGL}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$ one has $\mathbf{Z}\equiv S_{A}\otimes Z_{B}$, where $S_{A}=\mathrm{diag}(1,i)$ and $Z_{B}=\mathrm{diag}(1,-1)$.
  2. (ii)
    $g_{v}$ is entangling and CNOT-class: in $\operatorname{PGL}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$,
    $\mathbf{X}^{-1}\equiv\mathrm{CNOT}_{A,B}\,(X_{A}\otimes I_{B})$
    (projective equality), hence $g_{v}$ is locally equivalent to $\mathrm{CNOT}$ (Controlled-NOT).
  3. (iii)
    Consequently, if one can implement a universal single-qubit gate set on each factor and the monodromy gate $g_{v}$, then these gates generate a dense subgroup of $\operatorname{PU}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})\cong\operatorname{PU}(4)$.
Proof.
Along $\gamma_{u}$ the branch change $u^{1/4}\mapsto\zeta\,u^{1/4}$ is absorbed projectively by conjugation using $\mathbf{ZXZ}^{-1}=\zeta\mathbf{X}$, giving $g_{u}=[\mathbf{Z}]$. Similarly, along $\gamma_{v}$, one gets $g_{v}=[\mathbf{X}^{-1}]$. Under $|a,b\rangle\leftrightarrow|a+2b\rangle$,
$\mathbf{Z}|a,b\rangle=i^{a+2b}|a,b\rangle=i^{a}(-1)^{b}|a,b\rangle,$
so $\mathbf{Z}\equiv S_{A}\otimes Z_{B}$ in $\operatorname{PGL}$. A direct basis check shows $\mathbf{X}^{-1}$ agrees with $\mathrm{CNOT}_{A,B}(X_{A}\otimes I_{B})$ in $\operatorname{PGL}$, hence $g_{v}$ is locally equivalent to $\mathrm{CNOT}$. Finally, it is standard that a get set consisting of $\mathrm{CNOT}$ and universal single-qubit gates is universal on two qubits (up to phase), and local equivalence is implemented by single-qubit gates, proving (iii). ∎
Proposition 4.11 (Monodromy and failure of a Segre reduction).
Let $P\to X$ be a $\operatorname{PGL}_{n}$–torsor and fix $\mathbf{d}$ with stabilizer $G_{\mathbf{d}}\subset\operatorname{PGL}_{n}$. Suppose $U\to X$ is an fppf cover on which $P$ is trivial, so that the associated bundle $SB(\mathcal{A})|_{U}\simeq\mathbb{P}^{n-1}\times U$ carries the standard Segre subscheme $\Sigma_{\mathbf{d}}\times U$.
Then the following are equivalent:
[…]
Thus, a transition element $g\in\operatorname{PGL}_{n}$ is “non-entangling” (for type $\mathbf{d}$) precisely when $g(\Sigma_{\mathbf{d}})=\Sigma_{\mathbf{d}}$. In particular, if some transition element lies outside $G_{\mathbf{d}}$, then it sends some point of $\Sigma_{\mathbf{d}}$ outside $\Sigma_{\mathbf{d}}$.
Proof.
This is immediate from the definition of reduction of structure group and of the stabilizer $G_{\mathbf{d}}$. ∎
Remark 4.12 (Entangling holonomic gate).
From the perspective of quantum computing, Proposition 4.11 can be interpreted as follows. Choose an open set $U\subset X$ over which a $\mathbf{d}$-subsystem structure is trivialised, so that product states form the standard product-state locus $\Sigma_{\mathbf{d}}\subset\mathbb{P}^{n-1}$ on each fiber. For a loop $\gamma$ based in $U$, let $\rho(\gamma)\in\operatorname{PGL}_{n}$ denote the associated projective monodromy. Then:
$\rho(\gamma)\in G_{\mathbf{d}}\quad\Longleftrightarrow\quad\rho(\gamma)(\Sigma_{\mathbf{d}})=\Sigma_{\mathbf{d}}.$
In particular, if $\rho(\gamma)\notin G_{\mathbf{d}}$, then $\gamma$ acts as an entangling holonomic gate: there exists a product state $\psi]\in\Sigma_{\mathbf{d}}$ such that $\rho(\gamma)[\psi]\notin\Sigma_{\mathbf{d}}$.
Remark 4.13 (Brauer obstruction as the geometric source of entangling gates).
In the Kummer-symbol setting on $\mathcal{X}=(\mathbb{C}^{\times})^{2}$, the Brauer class of the degree-$4$ symbol algebra has period $4$, while a global $(2,2)$ subsystem structure would force the class to be $2$-torsion. Consequently no global reduction to $G_{(2,2)}$ exists, and any locally chosen tensor decomposition must acquire nonlocal monodromy around some loop. Theorem 4.10 makes this mechanism explicit: the $v$-loop produces a monodromy in the $\mathrm{CNOT}$-type, hence an entangling operation relative to the local $(2,2)$ splitting. From the Hilbert-scheme viewpoint, this monodromy is exactly the monodromy of the point in the subsystem-structure locus (Definition 2.18) rather than a symmetry inside the stabilizer.
Example 4.14 (Twisted $8$-level system on $X=\mathbb{G}_{\mathrm{m}}^{2}$ with locally multiple factorizations).
Let
$X:=\mathbb{G}_{\mathrm{m}}^{2}=\operatorname{Spec}\mathbb{C}[u^{\pm 1},v^{\pm 1}],\qquad n=8,$
so each fiber of the pure state space is (locally) $\mathbb{P}^{7}$.
Fix a primitive $8$th root $\zeta:=e^{2\pi i/8}$ and consider the $\mathcal{O}_{X}$-algebra
$A:=A(u,v)_{8}:=\mathcal{O}_{X}\langle x,y\rangle\Big/\big(x^{8}=u,\;y^{8}=v,\;yx=\zeta\,xy\big).$
Let $\pi:SB(A)\to X$ be the associated Severi–Brauer scheme (family of pure state spaces).
On a simply connected analytic open $U\subset X$ choose branches $u^{1/8},v^{1/8}$. Let $\mathcal{H}:=\mathbb{C}^{8}$ with basis $\{|r\rangle\}_{r\in\mathbb{Z}/8\mathbb{Z}}$ and define $\mathbf{X}$ and $\mathbf{Z}$ as eq. (4.1). Then
$x(u,v):=u^{1/8}\mathbf{X},\qquad y(u,v):=v^{1/8}\mathbf{Z}$
gives a local representation, hence
$A|_{U}\simeq\operatorname{End}(\mathcal{H}),\qquad SB(A)|_{U}\simeq\mathbb{P}(\mathcal{H})\times U.$
On $U$ one may identify $\mathcal{H}\simeq\mathbb{C}^{8}$ with either
$\text{(A) }\mathcal{H}\simeq\mathbb{C}^{2}\otimes\mathbb{C}^{2}\otimes\mathbb{C}^{2}\quad(d_{3}=(2,2,2)),$
via binary expansion $r=a+2b+4c$ $(a,b,c\in\{0,1\})$, i.e. $|a\rangle_{A}\otimes|b\rangle_{B}\otimes|c\rangle_{C}\leftrightarrow|r\rangle$, or
$\text{(B) }\mathcal{H}\simeq\mathbb{C}^{4}\otimes\mathbb{C}^{2}\quad(d_{2}=(4,2)),$
via $r=a+4b$ $(a\in\{0,1,2,3\},\,b\in\{0,1\})$, i.e. $|a\rangle_{A}\otimes|b\rangle_{B}\leftrightarrow|r\rangle$.
Let the fundamental loops be
$\gamma_{u}(t)=(u_{0}e^{2\pi it},v_{0}),\qquad\gamma_{v}(t)=(u_{0},v_{0}e^{2\pi it}).$
Analytic continuation multiplies $u^{1/8}$ or $v^{1/8}$ by $\zeta$. Compensating by conjugation yields projective monodromies
$g_{u}=[\mathbf{Z}]\in\operatorname{PGL}(\mathcal{H}),\qquad g_{v}=[\mathbf{X}^{-1}]\in\operatorname{PGL}(\mathcal{H}).$
Here it is important that $g_{v}$ is not a local operation: it entangles product states. To see this, we consider the following two cases (a) and (b).
(a) Under the $3$-qubit factorization $d_{3}=(2,2,2)$.  Take a local product state
$|\psi\rangle=(|0\rangle_{A}+|1\rangle_{A})\otimes|0\rangle_{B}\otimes|0\rangle_{C}\;\;\leftrightarrow\;\;|0\rangle+|1\rangle.$
Since $\mathbf{X}^{-1}|0\rangle=|7\rangle$ and $\mathbf{X}^{-1}|1\rangle=|0\rangle$,
$g_{v}|\psi\rangle\sim|7\rangle+|0\rangle\;\;\leftrightarrow\;\;|111\rangle+|000\rangle,$
which is the GHZ state (entangled).
(b) Under the bipartite factorization $d_{2}=(4,2)$.  Take
$|\phi\rangle=(|0\rangle_{A}+|1\rangle_{A})\otimes|0\rangle_{B}\;\;\leftrightarrow\;\;|0\rangle+|1\rangle.$
Then again $g_{v}|\phi\rangle\sim|7\rangle+|0\rangle$ corresponds to
$|0,0\rangle+|3,1\rangle\in\mathbb{C}^{4}\otimes\mathbb{C}^{2},$
which has Schmidt rank $\geq 2$, hence is entangled.
We now turn to the Brauer-period obstruction to global subsystem structures.  Let $\beta:=[A]\in\operatorname{Br}(X)$. By Lemma 4.5 we have $\operatorname{per}(\beta)=8$. For a subsystem type $d=(d_{1},\dots,d_{s})$ set $\ell=\mathrm{lcm}(d_{1},\dots,d_{s})$. Then
$d_{3}=(2,2,2)\Rightarrow\ell=2,\qquad d_{2}=(4,2)\Rightarrow\ell=4,$
so $\operatorname{per}(\beta)=8\nmid\ell$ in both cases. Hence, by Proposition 4.1, $\beta\notin\operatorname{Br}_{d_{3}}(X)$ and $\beta\notin\operatorname{Br}_{d_{2}}(X)$. By Theorem 2.10, the associated $\operatorname{PGL}_{8}$-torsor admits no reduction to $G_{d_{3}}$ nor $G_{d_{2}}$. Therefore the subsystem-structure loci $\operatorname{Hilb}^{\Sigma_{d_{i}}}(SB(A)/X)\simeq P/G_{d_{i}}$ admit no section $X\to P/G_{d_{i}}$ for $i=2,3$.
4.3. A twisted but $(p,p)$-reducible symbol tensor product on $\mathbb{G}_{\mathrm{m}}^{2}$
Throughout this subsection let $p$ be a prime, assume $\mathrm{char}(k)\nmid p$ and $k$ contains a primitive $p$-th root of unity $\zeta_{p}$. Set
$X=\mathbb{G}_{\mathrm{m}}^{2}=\operatorname{Spec}k[u^{\pm 1},v^{\pm 1}].$
Definition 4.15.
Fix $a,b\in k^{\times}$. Let $A_{u}$ (resp. $A_{v}$) be the sheaf of $\mathcal{O}_{X}$-algebras generated by $x_{u},y_{u}$ (resp. $x_{v},y_{v}$) with relations
$x_{u}^{p}=u,\quad y_{u}^{p}=a,\quad y_{u}x_{u}=\zeta_{p}x_{u}y_{u},\qquad x_{v}^{p}=v,\quad y_{v}^{p}=b,\quad y_{v}x_{v}=\zeta_{p}x_{v}y_{v}.$
Define the degree-$p^{2}$ Azumaya algebra
$A:=A_{u}\otimes_{\mathcal{O}_{X}}A_{v}.$
Proposition 4.16.
The algebra $A$ is Azumaya of degree $p^{2}$ on $X$. Its Brauer class $\beta:=[A]\in\operatorname{Br}(X)$ satisfies $\beta\in\operatorname{Br}(X)[p]$. Moreover, $A$ admits a global $(p,p)$-subsystem structure. Hence its Brauer class $\beta$ lies in $\operatorname{Br}_{(p,p)}(X)$.
Proof.
Azumaya-ness is checked after the Kummer cover adjoining $p$-th roots of $u$ and $v$: over $X^{\prime}=\operatorname{Spec}k[u^{\pm 1/p},v^{\pm 1/p}]$ one has $A_{u}|_{X^{\prime}}\simeq\operatorname{Mat}_{p}(\mathcal{O}_{X^{\prime}})$ and $A_{v}|_{X^{\prime}}\simeq\operatorname{Mat}_{p}(\mathcal{O}_{X^{\prime}})$, hence $A|_{X^{\prime}}\simeq\operatorname{Mat}_{p^{2}}(\mathcal{O}_{X^{\prime}})$. Thus $A$ is Azumaya of degree $p^{2}$.
Since $A]=[A_{u}]+[A_{v}]$ in $\operatorname{Br}(X)$ and each of $A_{u}],[A_{v}]$ is $p$-torsion, we have $\beta\in\operatorname{Br}(X)[p]$.
For reducibility, let $P_{u}\to X$ and $P_{v}\to X$ be the $\operatorname{PGL}_{p}$-torsors associated to $A_{u}$ and $A_{v}$. Then $P_{u}\times_{X}P_{v}$ is a $\operatorname{PGL}_{p}\times\operatorname{PGL}_{p}$-torsor. Pushing it forward along the tensor representation $\operatorname{PGL}_{p}\times\operatorname{PGL}_{p}\to\operatorname{PGL}_{p^{2}}$ yields the $\operatorname{PGL}_{p^{2}}$-torsor associated to $A$. Since the image lies in the stabilizer $G_{(p,p)}\subset\operatorname{PGL}_{p^{2}}$, this exhibits a $G_{(p,p)}$-reduction of $P(A)$, hence a $(p,p)$-subsystem structure. ∎
Remark 4.17.
Compare with the Kummer $p^{2}$-symbol class on the same base $X$: there one has $\mathrm{per}=p^{2}$, hence the torsion obstruction rules out membership in $\operatorname{Br}_{(p,p)}(X)$, although $SB(A)\to X$ exists. Proposition 4.16 shows that on $X=\mathbb{G}_{\mathrm{m}}^{2}$ there are also nontrivial twisted classes of degree $p^{2}$ that do globalize a $(p,p)$-subsystem.
4.4. A concrete quantum information system and a non-entangling monodromy
Fix a prime $p$ and put
$\mathcal{X}=(\mathbb{C}^{\times})^{2},\qquad(u,v)\in\mathcal{X},$
and let
$\zeta:=e^{2\pi i/p}.$
Fix nonzero constants $a,b\in\mathbb{C}^{\times}$ and choose once and for all $p$-th roots $\tilde{\alpha}:=a^{1/p}$ and $\tilde{\beta}:=b^{1/p}$. We describe an explicit local model realizing the symbol tensor product algebra of Definition 4.15 and show that the resulting gauge monodromies are non-entangling with respect to the bipartite type $(p,p)$.
Step 1: Local trivialization and a parameter-dependent representation. Choose a simply connected open set $U\subset\mathcal{X}$ on which we can pick analytic branches of the $p$-th roots $u^{1/p}$ and $v^{1/p}$. Fix Hilbert spaces
$\mathcal{H}_{A}:=\mathbb{C}^{p},\qquad\mathcal{H}_{B}:=\mathbb{C}^{p},\qquad\mathcal{H}:=\mathcal{H}_{A}\otimes\mathcal{H}_{B},$
with computational bases $\{|r\rangle_{A}\}_{r\in\mathbb{Z}/p\mathbb{Z}}$ and $\{|s\rangle_{B}\}_{s\in\mathbb{Z}/p\mathbb{Z}}$.
Define Weyl (shift/clock) operators on each factor by
$\mathbf{X}_{A}|r\rangle_{A}=|r+1\rangle_{A},\qquad\mathbf{Z}_{A}|r\rangle_{A}=\zeta^{\,r}|r\rangle_{A},$
$\mathbf{X}_{B}|s\rangle_{B}=|s+1\rangle_{B},\qquad\mathbf{Z}_{B}|s\rangle_{B}=\zeta^{\,s}|s\rangle_{B}.$
Then $\mathbf{Z}_{A}\mathbf{X}_{A}=\zeta\,\mathbf{X}_{A}\mathbf{Z}_{A}$ and $\mathbf{Z}_{B}\mathbf{X}_{B}=\zeta\,\mathbf{X}_{B}\mathbf{Z}_{B}$.
For $(u,v)\in U$, define operators on $\mathcal{H}$ by
$x_{u}(u,v):=u^{1/p}\,(\mathbf{X}_{A}\otimes\mathrm{id}_{\mathcal{H}_{B}}),\qquad y_{u}(u,v):=\tilde{\alpha}\,(\mathbf{Z}_{A}\otimes\mathrm{id}_{\mathcal{H}_{B}}),$
$x_{v}(u,v):=v^{1/p}\,(\mathrm{id}_{\mathcal{H}_{A}}\otimes\mathbf{X}_{B}),\qquad y_{v}(u,v):=\tilde{\beta}\,(\mathrm{id}_{\mathcal{H}_{A}}\otimes\mathbf{Z}_{B}).$
A direct computation gives
$x_{u}(u,v)^{p}=u\,\mathrm{id}_{\mathcal{H}},\qquad y_{u}(u,v)^{p}=a\,\mathrm{id}_{\mathcal{H}},\qquad y_{u}(u,v)\,x_{u}(u,v)=\zeta\,x_{u}(u,v)\,y_{u}(u,v),$
and similarly
$x_{v}(u,v)^{p}=v\,\mathrm{id}_{\mathcal{H}},\qquad y_{v}(u,v)^{p}=b\,\mathrm{id}_{\mathcal{H}},\qquad y_{v}(u,v)\,x_{v}(u,v)=\zeta\,x_{v}(u,v)\,y_{v}(u,v).$
Moreover, the $u$-generators commute with the $v$-generators since they act on different tensor factors. Thus, on $U$, the generators of
$A=A_{u}\otimes_{\mathcal{O}_{\mathcal{X}}}A_{v}$
from Definition 4.15 are realized concretely as the $(u,v)$-dependent operators $\{x_{u},y_{u},x_{v},y_{v}\}\subset\operatorname{End}(\mathcal{H})$.
Step 2: Monodromy along the basic loops. Let $(u_{0},v_{0})\in U$ be a base point and consider the loops
$\gamma_{u}(t)=(u_{0}e^{2\pi it},\,v_{0}),\qquad\gamma_{v}(t)=(u_{0},\,v_{0}e^{2\pi it})\qquad(0\leq t\leq 1).$
Analytic continuation of the chosen branches yields
$u^{1/p}\longmapsto\zeta\,u^{1/p}\quad\text{along }\gamma_{u},\qquad v^{1/p}\longmapsto\zeta\,v^{1/p}\quad\text{along }\gamma_{v}.$
Hence $x_{u}(u,v)$ picks up a factor $\zeta$ along $\gamma_{u}$, and $x_{v}(u,v)$ picks up a factor $\zeta$ along $\gamma_{v}$. These phase changes are absorbed by conjugation:
$y_{u}\,x_{u}\,y_{u}^{-1}=\zeta\,x_{u},\qquad y_{v}\,x_{v}\,y_{v}^{-1}=\zeta\,x_{v}.$
Therefore the corresponding projective monodromies can be taken as
$g_{u}=[\,y_{u}(u_{0},v_{0})\,]\in\operatorname{PGL}(\mathcal{H}),\qquad g_{v}=[\,y_{v}(u_{0},v_{0})\,]\in\operatorname{PGL}(\mathcal{H}).$
(Overall scalars are physically invisible on pure states, hence we work in $\operatorname{PGL}(\mathcal{H})$.) Since $y_{u}$ and $y_{v}$ act on different tensor factors, their lifts already commute in $\operatorname{GL}(\mathcal{H})$:
$g_{u}g_{v}g_{u}^{-1}g_{v}^{-1}\ \sim\ (\mathbf{Z}_{A}\otimes\mathrm{id})(\mathrm{id}\otimes\mathbf{Z}_{B})(\mathbf{Z}_{A}^{-1}\otimes\mathrm{id})(\mathrm{id}\otimes\mathbf{Z}_{B}^{-1})=\mathrm{id}_{\mathcal{H}}.$
Thus (in contrast to the Kummer $p^{2}$-symbol case) there is no entangling cross-effect in these basic monodromies.
Step 3: Non-entangling nature with respect to the bipartite type $(p,p)$. In the tensor-product picture $\mathcal{H}=\mathcal{H}_{A}\otimes\mathcal{H}_{B}$, the monodromies are explicitly local:
$g_{u}\sim\mathbf{Z}_{A}\otimes\mathrm{id}_{\mathcal{H}_{B}},\qquad g_{v}\sim\mathrm{id}_{\mathcal{H}_{A}}\otimes\mathbf{Z}_{B}.$
Hence $g_{u}$ and $g_{v}$ lie in the local-operation subgroup $\operatorname{PGL}(\mathcal{H}_{A})\times\operatorname{PGL}(\mathcal{H}_{B})\subset\operatorname{PGL}(\mathcal{H})$, so they preserve the product-state locus and cannot generate entanglement from a product state.
For instance, consider the product state
$|\psi\rangle:=(|0\rangle_{A}+|1\rangle_{A})\otimes|0\rangle_{B}.$
Applying $g_{u}$ gives
$g_{u}|\psi\rangle\ \sim\ (\mathbf{Z}_{A}\otimes\mathrm{id})\bigl((|0\rangle_{A}+|1\rangle_{A})\otimes|0\rangle_{B}\bigr)=(|0\rangle_{A}+\zeta\,|1\rangle_{A})\otimes|0\rangle_{B},$
which is still a product state (Schmidt rank $1$). Similarly,
$g_{v}|\psi\rangle\ \sim\ (|0\rangle_{A}+|1\rangle_{A})\otimes(\mathbf{Z}_{B}|0\rangle_{B})=(|0\rangle_{A}+|1\rangle_{A})\otimes|0\rangle_{B}$
(up to a phase), again a product state. Thus the gauge monodromy in this model is non-entangling with respect to $(p,p)$.
Remark 4.18 (The case $p=2$).
When $p=2$, one has $\mathcal{H}=\mathbb{C}^{2}\otimes\mathbb{C}^{2}$ and the monodromies are
$g_{u}\sim\sigma_{z}\otimes\mathrm{id},\qquad g_{v}\sim\mathrm{id}\otimes\sigma_{z}$
with Pauli $\sigma_{z},\sigma_{x}$, so they are manifestly local phase flips and (in contrast to the example in §4.2) never produce Bell-type entanglement.
4.5. Period is not sufficient: indecomposable classes of exponent $p$ and index $p^{2}$
Lemma 4.19.
Let $F$ be a field and $d=(p,p)$. If $\beta\in\operatorname{Br}_{(p,p)}(F)$, then there exists a separable extension $L/F$ of degree dividing $2$ such that $\beta_{L}$ is represented by a tensor product of two degree-$p$ central simple $L$-algebras.
Proof.
By definition, $\beta$ comes from a $G_{(p,p)}$-reduction of a $\operatorname{PGL}_{p^{2}}$-torsor. When $d=(p,p)$, the stabilizer has a component group $\pi_{0}(G_{(p,p)})\simeq\mathfrak{S}_{2}$, so the reduction determines (and is determined by) a possibly nontrivial $\mathfrak{S}_{2}$-torsor. After base change to its splitting field $L/F$ (degree $\leq 2$), the reduction lands in the identity component, which is the image of $\operatorname{PGL}_{p}\times\operatorname{PGL}_{p}$ under the tensor representation. Over $L$, this yields an ordered bipartite reduction, hence a decomposition of the associated degree-$p^{2}$ central simple algebra as a tensor product of two degree-$p$ algebras. ∎
Theorem 4.20.
Let $p$ be an odd prime. There exists a field $F$ of characteristic 0 and a Brauer class $\beta\in\operatorname{Br}(F)$ such that
$\exp(\beta)=p,\qquad\mathrm{ind}(\beta)=p^{2},$
and $\beta$ remains indecomposable after any prime-to-$p$ field extension. In particular, $\beta\notin\operatorname{Br}_{(p,p)}(F)$ although $\beta\in\operatorname{Br}(F)[p]$.
Proof.
The existence of such $\beta$ (indecomposable of exponent $p$ and index $p^{2}$, and stable under prime-to-$p$ extensions) is provided by the theory of generic abelian crossed products (see [McKinniePrimeToP2007, Theorem 3.16]). Assuming $\beta\in\operatorname{Br}_{(p,p)}(F)$, Lemma 4.19 would imply that $\beta$ becomes decomposable after a quadratic extension of $F$, which is a prime-to-$p$ extension. This contradicts the prime-to-$p$ indecomposability. ∎
Remark 4.21.
The torsion obstruction $\operatorname{Br}_{(p,p)}(F)\subset\operatorname{Br}(F)[p]$ is therefore sharp but not sufficient: there exist classes of exponent $p$ which cannot arise from any global $(p,p)$-subsystem structure.
4.6. A numerical split obstruction on curves (including elliptic curves)
Let $C$ be a smooth projective curve over an algebraically closed field $k$. Since $\operatorname{Br}(C)=0$, every $\mathbb{P}^{n-1}$-bundle over $C$ is of the form $\mathbb{P}(V)$ for some rank-$n$ vector bundle $V$, unique up to twisting by a line bundle.
Definition 4.22.
For a $\mathbb{P}^{n-1}$-bundle $P\simeq\mathbb{P}(V)\to C$, define
$\deg(P)\in\mathbb{Z}/n\mathbb{Z},\qquad\deg(P):=\deg(V)\ \mathrm{mod}\ n.$
This is well-defined because $\deg(V\otimes L)=\deg(V)+n\deg(L)$.
Proposition 4.23.
Let $d=(d_{A},d_{B})$ with $n=d_{A}d_{B}$ and set $g:=\gcd(d_{A},d_{B})$. If a $\mathbb{P}^{n-1}$-bundle $P\to C$ admits a $d$-subsystem structure (equivalently, is $d$-reducible), then $\deg(P)\equiv 0\ (\mathrm{mod}\ g)$ in $\mathbb{Z}/g\mathbb{Z}$.
Proof.
Choose $V$ with $P\simeq\mathbb{P}(V)$. If $P$ is $d$-reducible, then (for $d_{A}\neq d_{B}$, or after an étale cover removing the permutation ambiguity when $d_{A}=d_{B}$) one has an isomorphism
$V\;\simeq\;(V_{A}\otimes V_{B})\otimes L$
for some rank-$d_{A}$ bundle $V_{A}$, rank-$d_{B}$ bundle $V_{B}$, and a line bundle $L$. Taking degrees gives
$\deg(V)\equiv\deg(V_{A}\otimes V_{B})\equiv d_{B}\deg(V_{A})+d_{A}\deg(V_{B})\equiv 0\pmod{g}$,
and twisting by $L$ does not change $\deg(V)\pmod{g}$ since $n\equiv 0\pmod{g}$. Thus $\deg(P)\equiv 0\pmod{g}$. ∎
Example 4.24.
Let $E$ be an elliptic curve, let $n=p^{2}$ and $d=(p,p)$ with $p$ prime. Pick a line bundle $M$ of degree $1$ and set $V:=\mathcal{O}_{E}^{\oplus(n-1)}\oplus M$. Then $\mathbb{P}(V)\to E$ has $\deg(\mathbb{P}(V))\equiv 1\ (\mathrm{mod}\ p)$, hence it is not $(p,p)$-reducible by Proposition 4.23, although it is split (Brauer class 0).
4.7. Azumaya algebras from moduli of vector bundles
Let $C$ be a smooth projective curve over an algebraically closed field $k$. Fix integers $r\geq 2$ and a line bundle $\mathcal{L}$ on $C$ of degree $d$. Let $M:=M_{C}(r,\mathcal{L})$ denote the moduli space of stable vector bundles on $C$ of rank $r$ with fixed determinant $\mathcal{L}$.
Theorem 4.25 (Biswas–Sengupta [2018arXiv180402494B]).
The Brauer group $\operatorname{Br}(M)$ is cyclic of order $\gcd(r,d)$. Moreover, the Brauer class of the universal projective bundle restricted to $\{x\}\times M$ (for any point $x\in C$) generates $\operatorname{Br}(M)$.
Corollary 4.26.
Let $r=p^{2}$ with $p$ prime, and assume $\gcd(r,d)=p^{2}$ (e.g. $d\equiv 0\ (\mathrm{mod}\ p^{2})$). Let $\alpha\in\operatorname{Br}(M)$ be the generator from Theorem 4.25. Then $\operatorname{per}(\alpha)=p^{2}$, hence for the bipartite type $(p,p)$ one has
$\alpha\notin\operatorname{Br}_{(p,p)}(M).$
Equivalently, although the Severi–Brauer fibration $SB(A_{\alpha})\to M$ exists, the $(p,p)$-subsystem structure cannot be globalized on $M$.
Proof.
By Theorem 4.25, $\operatorname{per}(\alpha)=\gcd(r,d)=p^{2}$. For $d=(p,p)$ one has $\ell=\mathrm{lcm}(p,p)=p$, hence $\operatorname{Br}_{(p,p)}(M)\subset\operatorname{Br}(M)[p]$ by the torsion obstruction. Therefore $\alpha\notin\operatorname{Br}_{(p,p)}(M)$. ∎
4.8. Non-Brauer invariance and split classification on $\mathbb{P}^{1}$
4.8.1. Same Brauer class, different subsystem behavior
It is important to note that subsystem reducibility is a property of the $\operatorname{PGL}_{n}$–torsor, not only of the Brauer class. We record this explicitly.
Proposition 4.27 (Subsystem reducibility is not a Brauer invariant).
Fix a subsystem type $\mathbf{d}=(d_{1},\dots,d_{s})$ with $n=\prod_{i}d_{i}$ and assume that $\Sigma_{\mathbf{d}}\subset\mathbb{P}^{n-1}$ is a proper subscheme (thus, at least two of the $d_{i}$ are >1>1). Then there exist a scheme $X$ and Azumaya algebras $\mathcal{A},\mathcal{A}^{\prime}$ of degree $n$ on $X$ such that
$\mathcal{A}]=[\mathcal{A}^{\prime}]\in\operatorname{Br}(X),$
but $\mathcal{A}$ admits a subsystem structure of type $\mathbf{d}$ whereas $\mathcal{A}^{\prime}$ does not. In fact, one may take $X=\mathbb{P}^{1}_{k}$ and $\mathcal{A}]=[\mathcal{A}^{\prime}]=0$.
Proof.
Take $X=\mathbb{P}^{1}_{k}$, so $\operatorname{Br}(X)=0$.
Step 1: A reducible representative. Let
$E_{\mathrm{red}}:=\mathcal{O}_{\mathbb{P}^{1}}^{\oplus n}.$
Then $\mathcal{A}:=\underline{\operatorname{End}}(E_{\mathrm{red}})$ is split, hence $\mathcal{A}]=0$. Moreover $E_{\mathrm{red}}$ admits the evident tensor factorization
$E_{\mathrm{red}}\ \cong\ \bigl(\mathcal{O}^{\oplus d_{1}}\bigr)\otimes\cdots\otimes\bigl(\mathcal{O}^{\oplus d_{s}}\bigr),$
so the associated $\operatorname{PGL}_{n}$–torsor $\mathbb{P}(E_{\mathrm{red}})\to\mathbb{P}^{1}$ reduces to the stabilizer $G_{\mathbf{d}}$. Thus $\mathcal{A}$ admits a subsystem structure of type $\mathbf{d}$.
Step 2: A non-reducible representative with the same Brauer class. Let
$E_{\mathrm{irr}}:=\mathcal{O}_{\mathbb{P}^{1}}^{\oplus(n-1)}\oplus\mathcal{O}_{\mathbb{P}^{1}}(1),\qquad\mathcal{A}^{\prime}:=\underline{\operatorname{End}}(E_{\mathrm{irr}}).$
Again $\mathcal{A}^{\prime}$ is split, so $\mathcal{A}^{\prime}]=0=[\mathcal{A}]$.
We claim that $\mathcal{A}^{\prime}$ admits no globally compatible subsystem structure of type $\mathbf{d}$. Suppose to the contrary that the associated $\operatorname{PGL}_{n}$–torsor $\mathbb{P}(E_{\mathrm{irr}})\to\mathbb{P}^{1}$ admits a reduction to $G_{\mathbf{d}}$.
Since $k$ is algebraically closed, $\pi_{1}^{\mathrm{\acute{e}t}}(\mathbb{P}^{1}_{k})=1$, hence every finite étale torsor on $\mathbb{P}^{1}_{k}$ is trivial. In particular, any component-group (permutation) part of $G_{\mathbf{d}}$ produces no monodromy. Thus (after a global choice of ordering of equal factors) the reduction may be viewed as a reduction to the identity component, which is the image of $\prod_{i}\operatorname{PGL}_{d_{i}}$.
Since $\operatorname{Br}(\mathbb{P}^{1})=0$, each induced $\operatorname{PGL}_{d_{i}}$–torsor lifts to a $\operatorname{GL}_{d_{i}}$–torsor, hence to a rank-$d_{i}$ vector bundle $F_{i}$. The tensor product representation then gives an identification of $\operatorname{PGL}_{n}$–torsors
$\mathbb{P}(E_{\mathrm{irr}})\ \simeq\ \mathbb{P}(F_{1}\otimes\cdots\otimes F_{s}).$
By Lemma 4.29, this implies
$E_{\mathrm{irr}}\ \simeq\ (F_{1}\otimes\cdots\otimes F_{r})\otimes L$
for some line bundle $L$ on $\mathbb{P}^{1}$.
Now write the Grothendieck splitting of each $F_{i}$ as $F_{i}\simeq\bigoplus_{j=1}^{d_{i}}\mathcal{O}(a_{ij})$ with $a_{i1}\leq\cdots\leq a_{id_{i}}$. Let $m_{i}$ be the multiplicity of the minimal degree $a_{i1}$ inside $\{a_{ij}\}_{j}$. Then the tensor product splits as a direct sum of line bundles $\mathcal{O}\!\left(\sum_{i}a_{ij_{i}}\right)$, and the minimal degree occurs with multiplicity
[…]
Twisting by $L$ shifts all degrees uniformly and does not change this multiplicity.
On the other hand, $E_{\mathrm{irr}}=\mathcal{O}^{\oplus(n-1)}\oplus\mathcal{O}(1)$ has normalized splitting type $\{0,\dots,0,1\}$, so the minimal degree occurs with multiplicity n1n-1. Hence we must have $\prod_{i}m_{i}=n-1$.
But each $m_{i}$ satisfies $1\leq m_{i}\leq d_{i}$, and since $\Sigma_{\mathbf{d}}$ is proper we have at least two factors $d_{i}>1$, so for every $i$ we have
$\frac{n}{d_{i}}=\prod_{j\neq i}d_{j}\ \geq\ 2.$
If for some $i$ we had $m_{i}\leq d_{i}-1$, then
$\prod_{j=1}^{s}m_{j}\ \leq\ (d_{i}-1)\prod_{j\neq i}d_{j}\;=\;n-\frac{n}{d_{i}}\ \leq\ n-2,$
so it could not equal n1n-1. Therefore we must have $m_{i}=d_{i}$ for all $i$, and then $\prod_{i}m_{i}=n$, again impossible. This contradiction shows that no such decomposition of $E_{\mathrm{irr}}$ exists, hence $\mathbb{P}(E_{\mathrm{irr}})$ does not reduce to $G_{\mathbf{d}}$ and $\mathcal{A}^{\prime}$ admits no subsystem structure of type $\mathbf{d}$.
Thus $\mathcal{A}]=[\mathcal{A}^{\prime}]=0$ but only $\mathcal{A}$ is subsystem-reducible of type $\mathbf{d}$, as claimed. ∎
Remark 4.28 (Moral).
Even for $\beta=0$, subsystem structures are not automatic. Thus the entanglement geometry is extra datum beyond the Brauer class.
4.8.2. Split Brauer class on $\mathbb{P}^{1}$: classification of subsystem reducibility
Throughout this subsection let $k$ be algebraically closed of characteristic 0 and set $X=\mathbb{P}^{1}_{k}$. In particular $\operatorname{Br}(X)=0$.
Lemma 4.29 (Projective bundles on $\mathbb{P}^{1}$ and splitting types).
Let $P\to\mathbb{P}^{1}$ be a principal $\operatorname{PGL}_{n}$–torsor. Then there exists a rank-$n$ vector bundle $E$ on $\mathbb{P}^{1}$ such that $P\simeq\mathbb{P}(E)$. Moreover, if $E$ and $E^{\prime}$ are rank-$n$ bundles with $\mathbb{P}(E)\simeq\mathbb{P}(E^{\prime})$, then $E^{\prime}\simeq E\otimes L$ for some line bundle $L\in\operatorname{Pic}(\mathbb{P}^{1})$.
Consequently, after choosing any such lift $E$ and writing its Grothendieck splitting
$E\;\cong\;\bigoplus_{m=1}^{n}\mathcal{O}_{\mathbb{P}^{1}}(a_{m})\qquad(a_{1}\leq\cdots\leq a_{n}),$
the multiset $\{a_{m}-a_{1}\}_{m=1}^{n}$ depends only on the isomorphism class of $P$.
Proof.
Since $\operatorname{Br}(\mathbb{P}^{1})=0$, the connecting map $H^{1}(\mathbb{P}^{1},\operatorname{PGL}_{n})\to H^{2}(\mathbb{P}^{1},\mathbb{G}_{\mathrm{m}})=\operatorname{Br}(\mathbb{P}^{1})$ is trivial, so $H^{1}(\mathbb{P}^{1},\operatorname{GL}_{n})\to H^{1}(\mathbb{P}^{1},\operatorname{PGL}_{n})$ is surjective. Thus $P$ admits a lift to a $\operatorname{GL}_{n}$–torsor, so $P\simeq\mathbb{P}(E)$ for some vector bundle $E$.
For the uniqueness statement, apply nonabelian cohomology to $1\to\mathbb{G}_{\mathrm{m}}\to\operatorname{GL}_{n}\to\operatorname{PGL}_{n}\to 1$. The kernel of $H^{1}(\mathbb{P}^{1},\operatorname{GL}_{n})\to H^{1}(\mathbb{P}^{1},\operatorname{PGL}_{n})$ is the image of $H^{1}(\mathbb{P}^{1},\mathbb{G}_{\mathrm{m}})=\operatorname{Pic}(\mathbb{P}^{1})$. Concretely this says that two lifts differ by tensoring with a line bundle.
Finally, Grothendieck splitting holds for any vector bundle on $\mathbb{P}^{1}$, and twisting by a line bundle shifts all $a_{m}$ by the same integer, so the normalized multiset $\{a_{m}-a_{1}\}$ is intrinsic to $P$. ∎
Theorem 4.30 (A splitting-type criterion for bipartite tensor reducibility on $\mathbb{P}^{1}$).
Fix a bipartite type $\mathbf{d}=(d_{A},d_{B})$ and set $n=d_{A}d_{B}$. Let $P\to\mathbb{P}^{1}$ be a principal $\operatorname{PGL}_{n}$–torsor, and choose a lift $P\simeq\mathbb{P}(E)$ as in Lemma 4.29. Write
$E\;\cong\;\bigoplus_{m=1}^{n}\mathcal{O}_{\mathbb{P}^{1}}(a_{m})\qquad(a_{1}\leq\cdots\leq a_{n}).$
Then the following are equivalent:
  1. (1)
    $P$ is $G_{\mathbf{d}}$–reducible (admits a reduction of structure group to the stabilizer $G_{\mathbf{d}}\subset\operatorname{PGL}_{n}$), so the split Azumaya algebra $\underline{\operatorname{End}}(E)$ admits a subsystem structure of type $\mathbf{d}$.
  2. (2)
    There exist vector bundles $F$ and $G$ on $\mathbb{P}^{1}$ of ranks $d_{A}$ and $d_{B}$ and a line bundle $L$ such that
    $E\;\cong\;(F\otimes G)\otimes L.$
  3. (3)
    (Additive decomposition of splitting type) There exist integers $b_{1}\leq\cdots\leq b_{d_{A}}$, $c_{1}\leq\cdots\leq c_{d_{B}}$ and an integer $t$ such that the multiset $\{a_{m}\}_{m=1}^{n}$ equals the multiset
    $\{\,b_{i}+c_{j}+t\mid 1\leq i\leq d_{A},\ 1\leq j\leq d_{B}\,\}.$
In particular, the $G_{\mathbf{d}}$–reducible locus inside the split Brauer fiber $\delta^{-1}(0)=H^{1}(\mathbb{P}^{1},\operatorname{PGL}_{n})$ is a proper subset (unless $(d_{A},d_{B})=(1,n)$ or $(n,1)$), described by the purely numerical condition (3).
Proof.
[…]
$(F\otimes G)\otimes L\;\simeq\;\bigoplus_{i,j}\mathcal{O}(b_{i}+c_{j}+t),$
and conversely any such multiset determines $F,G,L$.
[…]
$g^{E}_{\alpha\beta}\;=\;(g^{F}_{\alpha\beta}\otimes g^{G}_{\alpha\beta})\cdot\lambda_{\alpha\beta}\in\operatorname{GL}_{n},$
where $\lambda_{\alpha\beta}\in\mathbb{G}_{\mathrm{m}}$ is the scalar transition of $L$. Passing to $\operatorname{PGL}_{n}$ kills scalars, so the induced cocycle lies in the image of the tensor product subgroup, hence in $G_{\mathbf{d}}\subset\operatorname{PGL}_{n}$. This is exactly a reduction of $P$ to $G_{\mathbf{d}}$.
[…]
Thus (1) yields principal $\operatorname{PGL}_{d_{A}}$– and $\operatorname{PGL}_{d_{B}}$–torsors $Q_{A}$ and $Q_{B}$ on $\mathbb{P}^{1}$ whose induced $\operatorname{PGL}_{n}$–torsor (via the tensor representation) is isomorphic to $P$. Because $\operatorname{Br}(\mathbb{P}^{1})=0$, both $Q_{A}$ and $Q_{B}$ lift to vector bundles $F$ and $G$ of ranks $d_{A}$ and $d_{B}$. The induced $\operatorname{PGL}_{n}$–torsor is then $\mathbb{P}(F\otimes G)$ by inspection of transition functions. Hence $\mathbb{P}(E)\simeq\mathbb{P}(F\otimes G)$, and Lemma 4.29 implies $E\simeq(F\otimes G)\otimes L$ for some line bundle $L$. ∎
Corollary 4.31 (Complete criterion for $(2,2)$ on $\mathbb{P}^{1}$).
Let $n=4$ and $\mathbf{d}=(2,2)$. Let $P\simeq\mathbb{P}(E)\to\mathbb{P}^{1}$ with $E\simeq\mathcal{O}(a_{1})\oplus\mathcal{O}(a_{2})\oplus\mathcal{O}(a_{3})\oplus\mathcal{O}(a_{4})$ and $a_{1}\leq a_{2}\leq a_{3}\leq a_{4}$.
Then $P$ is $G_{(2,2)}$–reducible if and only if
$a_{1}+a_{4}\;=\;a_{2}+a_{3}.$
So, after twisting $E$ so that $a_{1}=0$, the normalized splitting type is
$(0,\ b,\ c,\ b+c)\qquad\text{for some integers }0\leq b\leq c.$
When the condition holds, one can take explicitly
$F=\mathcal{O}(a_{1})\oplus\mathcal{O}(a_{3}),\qquad G=\mathcal{O}(0)\oplus\mathcal{O}(a_{2}-a_{1}),$
so that $E\simeq F\otimes G$ and the $(2,2)$ subsystem structure is induced from this factorization.
Proof.
Apply Theorem 4.30 with $d_{A}=d_{B}=2$. In the $2\times 2$ case, an additive decomposition of the multiset $\{a_{m}\}$ is equivalent to the “parallelogram” condition $a_{1}+a_{4}=a_{2}+a_{3}$ (after ordering). The explicit $F,G$ are obtained by the constructive choice explained in the statement. ∎
Remark 4.32.
Even though $\beta=0\in\operatorname{Br}_{(2,2)}(\mathbb{P}^{1})$ (so some representative admits a subsystem structure), Corollary 4.31 shows that many split representatives do not. For example, the splitting types $(0,0,1,1)$ satisfy $0+1=0+1$ and are (2,2)(2,2)–reducible, whereas $(0,0,0,1)$ fail the condition and are not (2,2)(2,2)–reducible. Thus, already over $\mathbb{P}^{1}$ the existence of entanglement geometry is controlled by the finer $\operatorname{PGL}_{n}$–torsor class, not by the Brauer class alone.
Example 4.33 (An explicit split example on $\mathbb{P}^{1}$ for $\mathbf{d}=(2,2)$).
Let $X=\mathbb{P}^{1}_{k}$, $n=4$, and $\mathbf{d}=(2,2)$.
(i) (2,2)(2,2)–reducible. Take $E_{\mathrm{red}}:=\mathcal{O}^{\oplus 2}\oplus\mathcal{O}(1)^{\oplus 2}$, so that $a_{1}+a_{4}=a_{2}+a_{3}$ holds and $\mathbb{P}(E_{\mathrm{red}})$ is (2,2)(2,2)–reducible by Corollary 4.31.
(ii) not (2,2)(2,2)–reducible. Take $E_{\mathrm{irr}}:=\mathcal{O}^{\oplus 3}\oplus\mathcal{O}(1)$, for which $a_{1}+a_{4}\neq a_{2}+a_{3}$. Hence $\mathbb{P}(E_{\mathrm{irr}})$ is not (2,2)(2,2)–reducible by Corollary 4.31, even though both bundles define split Azumaya algebras (Brauer class 0).
5. Numerical invariants of Schmidt-rank loci
Throughout this section we are in the bipartite case $\mathbf{d}=(d_{A},d_{B})$ with $n=d_{A}d_{B}$. Let
$R_{\leq r}\subset\mathbb{P}(\,k^{d_{A}}\otimes k^{d_{B}}\,)$
denote the classical determinantal variety of tensors (matrices) of rank $\leq r$.
For each geometric point $x\to X$, choose an identification $SB(\mathcal{A})_{x}\simeq\mathbb{P}^{n-1}$ and let
$H_{x}\in\operatorname{Pic}\bigl(SB(\mathcal{A})_{x}\bigr)\cong\mathbb{Z}$
denote the resulting hyperplane class, namely the class of $\mathcal{O}_{\mathbb{P}^{n-1}}(1)$. This class is independent of the chosen identification, since any two such identifications differ by an element of $\operatorname{PGL}_{n}$ and the $\operatorname{PGL}_{n}$–action preserves $\mathcal{O}_{\mathbb{P}^{n-1}}(1)$.
Throughout Sec. 5, whenever we speak of the degree or the Hilbert polynomial of a fiber $(\Sigma_{\leq r})_{x}\subset SB(\mathcal{A})_{x}$, it is always taken with respect to the standard fiberwise polarization $\mathcal{O}_{SB(\mathcal{A})_{x}}(1)$ (equivalently, with respect to $H_{x}$) and its restriction to $(\Sigma_{\leq r})_{x}$. In the split model $R_{\leq r}\subset\mathbb{P}(V_{A}\otimes V_{B})$, this agrees with the standard polarization coming from $\mathcal{O}_{\mathbb{P}(V_{A}\otimes V_{B})}(1)$.
Proposition 5.1 (Basic dimension data).
For $1\leq r\leq\min(d_{A},d_{B})$ the determinantal variety $R_{\leq r}$ has
$\mathrm{codim}_{\mathbb{P}^{d_{A}d_{B}-1}}(R_{\leq r})=(d_{A}-r)(d_{B}-r),$
equivalently
$\dim(R_{\leq r})=r(d_{A}+d_{B}-r)-1.$
Proof.
Standard determinantal geometry: the affine cone is the rank $\leq r$ locus in $\operatorname{Mat}_{d_{A}\times d_{B}}$, which has codimension $(d_{A}-r)(d_{B}-r)$. Projectivising decreases dimension by $1$. ∎
Proposition 5.2 (Degree of the determinantal locus).
Assume $d_{A}\leq d_{B}$ for definiteness. The projective degree of $R_{\leq r}$ in $\mathbb{P}^{d_{A}d_{B}-1}$ is
$\deg(R_{\leq r})\;=\;\prod_{i=0}^{d_{A}-r-1}\frac{(d_{B}+i)!\,i!}{(r+i)!\,(d_{B}-r+i)!}.$
In particular, for $r=1$ (the product-state locus) one recovers $\deg(\mathbb{P}^{d_{A}-1}\times\mathbb{P}^{d_{B}-1})=\binom{d_{A}+d_{B}-2}{d_{A}-1}$.
Sketch of proof.
This follows from the Thom–Porteous formula applied to the universal flattening map on projective space, and by extracting the top coefficient against the hyperplane class. The displayed closed form is the standard simplification of that determinantal class. ∎
Let $S:=k[z_{ij}\mid 1\leq i\leq d_{A},\ 1\leq j\leq d_{B}]$ and let $I_{r+1}$ be the ideal of $(r+1)\times(r+1)$ minors. Then $R_{\leq r}=\operatorname{Proj}(S/I_{r+1})$.
Proposition 5.3 (Representation-theoretic Hilbert series).
The graded coordinate ring $S/I_{r+1}$ has the following decomposition as a $\operatorname{GL}_{d_{A}}\times\operatorname{GL}_{d_{B}}$–representation:
$(S/I_{r+1})_{t}\;\cong\;\bigoplus_{\begin{subarray}{c}\lambda\vdash t\\ \ell(\lambda)\leq r\end{subarray}}\mathbb{S}_{\lambda}(k^{d_{A}})\ \otimes\ \mathbb{S}_{\lambda}(k^{d_{B}}),$
where $\lambda$ ranges over partitions of $t$ with at most $r$ parts and $\mathbb{S}_{\lambda}$ denotes the Schur functor. Consequently, the Hilbert series is
$\mathrm{Hilb}_{R_{\leq r}}(z)\;=\;\sum_{\lambda:\,\ell(\lambda)\leq r}\dim\mathbb{S}_{\lambda}(k^{d_{A}})\cdot\dim\mathbb{S}_{\lambda}(k^{d_{B}})\ z^{|\lambda|}.$
In particular, the Hilbert polynomial of $R_{\leq r}$ is determined explicitly by $(d_{A},d_{B},r)$.
Remark 5.4.
For $r=1$ (product-state), one gets a closed formula:
$h^{0}(\Sigma_{A,B},\mathcal{O}(t))=\binom{t+d_{A}-1}{d_{A}-1}\binom{t+d_{B}-1}{d_{B}-1}$,
hence the Hilbert polynomial is the product of two binomials. For general $r$, the decomposition above gives an effective algorithm (via hook-length formula for $\dim\mathbb{S}_{\lambda}$) to compute the Hilbert polynomial.
Theorem 5.5 (Relative constancy of numerics for Schmidt-rank locus).
Let $(\mathcal{A},P_{\mathbf{d}})$ be an Azumaya algebra on $X$ with a chosen bipartite type $\mathbf{d}=(d_{A},d_{B})$. For each $r$, the Schmidt-rank $\leq r$ locus $\Sigma_{\leq r}(\mathcal{A},\mathbf{d})\subset SB(\mathcal{A})$ is flat over $X$, and for every $x\in X$ the fiber $(\Sigma_{\leq r})_{x}$ is isomorphic to the classical determinantal variety $R_{\leq r}$. Hence:
  • $\dim((\Sigma_{\leq r})_{x})$ and $\operatorname{codim}((\Sigma_{\leq r})_{x}\subset SB(\mathcal{A})_{x})$ are constant in $x$ and given by Proposition 5.1;
  • the projective degree of each fiber is constant and given by Proposition 5.2;
  • the Hilbert polynomial of each fiber is constant and given by Proposition 5.3.
Proof.
By construction, $\Sigma_{\leq r}(\mathcal{A},\mathbf{d})$ is the associated bundle $P_{\mathbf{d}}\times^{G_{\mathbf{d}}}R_{\leq r}$, hence is locally (fppf) the product $R_{\leq r}\times U$. Flatness and fiber identification follow. Numerical invariants are then fiberwise identical to those of $R_{\leq r}$. ∎
Appendix A Physics example: Quantum spin-chain on a torus
Here we present an example of a quantum spin-chain system on a torus whose ground state is locally a product state, yet becomes entangled after gluing along a coordinate due to a geometric obstruction. The example has the same logical structure as Example 4.14 in the main text:
  1. (1)
    Locally, a tensor decomposition (a product-state locus) can be taken.
  2. (2)
    However, as one follows a loop in the base space, the monodromy leaves $G_{\mathbf{d}}$.
  3. (3)
    Because of this, the decomposition cannot be transported globally. As a result, entanglement arises as an obstruction originating from gluing.
The spin-chain model studied below can be viewed as a physical reformulation of the abstract obstruction statements in the main text (Corollary 2.13, Proposition 4.11, Proposition 4.1).
A.1. Parameter space and the one-magnon Hilbert space
Let
$X:=(\mathbb{C}^{\times})^{2},\qquad(u,v)\in X.$
For a Hermitian (physical) family one may restrict to the real two-torus
$X_{\mathrm{phys}}:=(S^{1})^{2}\subset(\mathbb{C}^{\times})^{2},\qquad u=e^{i\theta_{u}},\ v=e^{i\theta_{v}}.$
We focus on the case $p=2$, hence $m=p^{2}=4$ and $\zeta:=e^{2\pi i/m}=i$.
Consider a periodic spin-$\tfrac{1}{2}$ chain with four sites $r\in\mathbb{Z}/4\mathbb{Z}$. Let $\sigma_{r}^{\pm}$ denote the raising/lowering operators at site $r$, and let
$n_{r}:=\frac{1-\sigma_{r}^{z}}{2}$
be the down-spin (magnon) number operator.
Let $|\!\uparrow\uparrow\uparrow\uparrow\rangle$ be the fully polarized state, and define the one-magnon basis
$|r\rangle:=\sigma_{r}^{-}|\!\uparrow\uparrow\uparrow\uparrow\rangle,\qquad r\in\mathbb{Z}/4\mathbb{Z}.$
The one-magnon subspace
$\mathcal{H}:=\mathrm{span}\{|0\rangle,|1\rangle,|2\rangle,|3\rangle\}\cong\mathbb{C}^{4}$
is preserved by any Hamiltonian conserving total magnon number.
On $\mathcal{H}$ consider the operators (4.1)
(A.1) ${\mathbf{X}}|r\rangle=|r+1\rangle,\qquad{\mathbf{Z}}|r\rangle=\zeta^{\,r}|r\rangle\quad(\zeta=i),$
so that $\mathbf{ZX}=\zeta\mathbf{XZ}$. (Physically, $\mathbf{X}$ is the lattice translation restricted to $\mathcal{H}$, while $\mathbf{Z}$ can be implemented by a site-dependent phase gate by controlling a local magnetic field.)
A.2. A local Hamiltonian and its explicit ground state
Fix constants $J>0$ and $\Delta>J$. On any simply connected chart $U\subset X^{\mathrm{an}}$ choose an analytic branch $u^{1/4}$. Define a local, strictly short-range spin Hamiltonian on the full spin Hilbert space by
(A.2) $H_{U}(u,v):=-J\Big(u^{-1/4}\,\sigma_{0}^{+}\sigma_{1}^{-}+u^{1/4}\,\sigma_{0}^{-}\sigma_{1}^{+}\Big)\;+\;\Delta\,(n_{2}+n_{3}).$
This Hamiltonian preserves magnon number. In particular it restricts to an operator on $\mathcal{H}$. In the basis $\{|0\rangle,|1\rangle,|2\rangle,|3\rangle\}$ one has
(A.3) $H_{U}\big|_{\mathcal{H}}=\begin{pmatrix}0&-Ju^{1/4}&0&0\\ -Ju^{-1/4}&0&0&0\\ 0&0&\Delta&0\\ 0&0&0&\Delta\end{pmatrix}$.
Proposition A.1 (Local ground state on a chart).
Assume $\Delta>J>0$ and $|u|=1$ (so $|u^{1/4}|=1$). Then $H_{U}|_{\mathcal{H}}$ has eigenvalues $J,+J,\Delta,\Delta$, and its unique ground state (energy J-J) is
(A.4) $|\mathrm{GS}\rangle_{U}=\frac{1}{\sqrt{2}}\Big(u^{1/4}|0\rangle+|1\rangle\Big).$
Moreover, since the fully polarized state has energy 0 and all states with magnons on sites $2,3$ cost at least $\Delta>J$, this is also the ground state of the full Hamiltonian $H_{U}$.
Proof.
The $2\times 2$ block on $\mathrm{span}\{|0\rangle,|1\rangle\}$ has eigenvalues $\pm J$. The remaining basis states $|2\rangle,|3\rangle$ have energy $\Delta$. Since $\Delta>J$, the lowest eigenvalue is J-J, and a direct eigenvector computation yields (A.4). The final statement follows by comparing energies in different magnon-number sectors. ∎
A.3. Gluing by projective conjugation and entangling transport
The Hamiltonian (A.2) is written in a local coordinate/branch. To define a globally twisted family (a family defined only up to fiberwise conjugation), we specify its gluing on overlaps.
Let $T$ denote the unitary translation operator on the spin chain, characterized by $T\sigma_{r}^{\pm}T^{-1}=\sigma_{r+1}^{\pm}$ (indices mod $4$). Its restriction to $\mathcal{H}$ is exactly the Weyl shift $X$ in (A.1). Define, on another simply connected chart $U^{\prime}\subset X^{\mathrm{an}}$, the Hamiltonian by conjugation:
(A.5) $H_{U^{\prime}}(u,v):=T^{-1}H_{U}(u,v)\,T.$
In spin operators this is still strictly local:
(A.6) $H_{U^{\prime}}(u,v)=-J\Big(u^{-1/4}\,\sigma_{3}^{+}\sigma_{0}^{-}+u^{1/4}\,\sigma_{3}^{-}\sigma_{0}^{+}\Big)\;+\;\Delta\,(n_{1}+n_{2}).$
By construction, on $U\cap U^{\prime}$ the two local descriptions are related by a gauge change (conjugation) in the projective group.
Proposition A.2 (Ground-state gluing).
On $U\cap U^{\prime}$, the (projective) ground state lines glue via
$\mathrm{GS}]_{U^{\prime}}=[\mathbf{X}^{-1}\mathrm{GS}]_{U}\ \in\ \mathbb{P}(\mathcal{H}),$
hence
(A.7) $|\mathrm{GS}\rangle_{U^{\prime}}=\mathbf{X}^{-1}|\mathrm{GS}\rangle_{U}=\frac{1}{\sqrt{2}}\Big(u^{1/4}|3\rangle+|0\rangle\Big).$
Proof.
Since $H_{U^{\prime}}=\mathbf{X}^{-1}H_{U}\mathbf{X}$ on $\mathcal{H}$, eigenlines are transported by $\mathbf{X}^{-1}$. Applying $\mathbf{X}^{-1}$ to (A.4) yields (A.7). ∎
A.4. Local bipartite splitting and emergence of entanglement
On each simply connected chart we may impose a local bipartite identification $\mathcal{H}\cong\mathbb{C}^{2}\otimes\mathbb{C}^{2}$ by writing
$r=a+2b,\qquad a,b\in\{0,1\},$
and declaring
$|a\rangle_{A}\otimes|b\rangle_{B}\ \longleftrightarrow\ |r=a+2b\rangle.$
Equivalently,
$|0\rangle=|00\rangle,\quad|1\rangle=|10\rangle,\quad|2\rangle=|01\rangle,\quad|3\rangle=|11\rangle.$
Below, the Schmidt rank is taken with respect to the local identification $\mathcal{H}\cong\mathbb{C}^{2}\otimes\mathbb{C}^{2}$ introduced above (not the physical tensor product $(\mathbb{C}^{2})^{\otimes 4}$ of the full spin chain).
Proposition A.3 (Product locally, entangled after gluing).
With respect to the above local splitting $\mathcal{H}\cong\mathbb{C}^{2}\otimes\mathbb{C}^{2}$, the local ground state $|\mathrm{GS}\rangle_{U}$ is a product state (Schmidt rank $1$), whereas the glued ground state $|\mathrm{GS}\rangle_{U^{\prime}}$ is entangled (Schmidt rank $2$). More explicitly,
(A.8) $|\mathrm{GS}\rangle_{U}$ $=\frac{1}{\sqrt{2}}\Big(u^{1/4}|00\rangle+|10\rangle\Big)=\Big(\tfrac{u^{1/4}|0\rangle_{A}+|1\rangle_{A}}{\sqrt{2}}\Big)\otimes|0\rangle_{B},$
(A.9) $|\mathrm{GS}\rangle_{U^{\prime}}$ $=\frac{1}{\sqrt{2}}\Big(|00\rangle+u^{1/4}|11\rangle\Big),$
and the latter has Schmidt rank $2$. In particular, if $u=1$ then $|\mathrm{GS}\rangle_{U^{\prime}}\propto|00\rangle+|11\rangle$ is a Bell state.
Proof.
The expression for $|\mathrm{GS}\rangle_{U}$ is immediate from (A.4). For $|\mathrm{GS}\rangle_{U^{\prime}}$, use (A.7) and the dictionary $|3\rangle=|11\rangle$, $|0\rangle=|00\rangle$. The coefficient matrix of $|\mathrm{GS}\rangle_{U^{\prime}}$ in the product basis is diagonal with two nonzero entries, hence has rank $2$, so the Schmidt rank is 22. ∎
A.5. Interpretation: an intrinsically entangling phase from twisted gluing
The key point is that the transition element transporting local trivialisations is projectively $g_{v}=[\mathbf{X}^{-1}]\in\mathrm{PGL}(\mathcal{H})$. Relative to the local bipartite type $(2,2)$, this element is not a local operation: it does not lie in the stabilizer $G_{(2,2)}\subset\mathrm{PGL}_{4}$. Consequently, even though one can choose a local product-state locus on each chart, these loci cannot glue to a global relative Segre subscheme. Therefore, no global $(2,2)$ subsystem structure exists, and the family is intrinsically entangled (Definition 2.8).
Finally, the underlying obstruction appears as a scalar discrepancy in $\mathrm{GL}(\mathcal{H})$: if one also considers the monodromy in the $u$-direction, one can take $g_{u}=[\mathbf{Z}]$. Then the commutator of lifts is
$\mathbf{ZX^{-1}Z^{-1}X}=\zeta^{-1}\,\mathrm{id}_{\mathcal{H}}=i^{-1}\,\mathrm{id}_{\mathcal{H}},$
a nontrivial scalar of order $4$. This is the local manifestation of the Brauer twisting responsible for the failure of a global Segre reduction of type $(2,2)$.
Remark A.4.
Obstruction classes also appear in familiar settings in physics. For example, in condensed matter theory, band topology studies how the occupied-band eigenspaces of a gapped Bloch Hamiltonian over the Brillouin torus $T^{d}$ assemble into a complex vector bundle (the Bloch bundle). The nontriviality of its $U(r)$-valued gluing data underlies the Berry connection/curvature, and characteristic invariants such as Chern numbers may be viewed as detecting the obstruction to choosing a globally consistent frame (or trivialization) of the occupied subbundle.
By contrast, the obstruction considered in this paper concerns the global compatibility of subsystem structure. We allow the background family of pure-state spaces to globalize only projectively, and we ask whether a prescribed tensor-product type $\mathbf{d}$ can be made globally coherent. This amounts to asking equivalently whether the associated $\operatorname{PGL}_{n}$-torsor admits a reduction to the Segre stabilizer $G_{\mathbf{d}}$. When such a reduction fails, transition functions necessarily exit the local operation group, so a state that is locally product can become entangled after gluing. In this sense, our obstruction classes are not Berry/Chern-type invariants. They arise from globalizing the subsystem decomposition itself rather than from globalizing eigenstates or eigenbundles.
Appendix B Terms and notations
Symbol Meaning
$X$ Base scheme (classical parameter space).
$\mathcal{A}$ Azumaya algebra on $X$ (family of $n$-level quantum systems).
$n$ Degree of $\mathcal{A}$.
$P\to X$ Principal $\operatorname{PGL}_{n}$-torsor associated to $\mathcal{A}$.
$SB(\mathcal{A})\to X$ Severi–Brauer scheme (pure-state fibration).
$\mathbf{d}=(d_{1},\dots,d_{s})$ Subsystem type with $n=\prod_{i=1}^{s}d_{i}$.
$\Sigma_{\mathbf{d}}$ Segre variety (product-state locus in the split model).
$G_{\mathbf{d}}\subset\operatorname{PGL}_{n}$ Stabilizer of $\Sigma_{\mathbf{d}}$.
$P_{\mathbf{d}}\to X$ A subsystem structure: reduction of $P$ to $G_{\mathbf{d}}$.
$\Sigma_{\mathbf{d}}(\mathcal{A})\subset SB(\mathcal{A})$ Descended Segre subfibration (product-state locus in families).
$\Sigma_{\leq r}(\mathcal{A},\mathbf{d})\subset SB(\mathcal{A})$ When $r>1$, Schmidt-rank $\leq r$ locus. In general, $\Sigma_{\leq r}$ contains both entangled and pure states. When $r=1$, product-state locus.
$\Sigma_{=r}:=\Sigma_{\leq r}\setminus\Sigma_{\leq r-1}$ When $r>1$, Schmidt-rank $r$ stratum. Every state in $\Sigma_{=r}$ is entangled for $r>1$. When $r=1$, product-state stratum. ($\Sigma_{=0}=\emptyset$)
$\mathrm{SR}_{\mathbf{d}}([\psi])$ Schmidt-rank entanglement measure defined via the filtration $\Sigma_{\leq r}$.
$\widetilde{\Sigma}_{\leq r}(\mathcal{A},\mathbf{d})$ Relative incidence resolution of $\Sigma_{\leq k}(\mathcal{A},\mathbf{d})$.
$\rho_{r}:\widetilde{\Sigma}_{\leq r}\to\Sigma_{\leq r}$ Canonical (projective birational) resolution map.
$\operatorname{Br}(X)$ Brauer group of $X$.
$\beta=[\mathcal{A}]\in\operatorname{Br}(X)$ Brauer class of $\mathcal{A}$.
$\operatorname{Br}_{\mathbf{d}}(X)\subset\operatorname{Br}(X)$ Subset of classes admitting a subsystem structure of type $\mathbf{d}$.
$\ell=\operatorname{lcm}(d_{1},\dots,d_{s})$ Torsion bound appearing in the period constraint for $\operatorname{Br}_{\mathbf{d}}(X)$.
$\operatorname{Hilb}^{\Sigma_{\mathbf{d}}}(SB(\mathcal{A})/X)$ subsystem-structure locus in the relative Hilbert scheme.
$P/G_{\mathbf{d}}(\simeq\operatorname{Hilb}^{\Sigma_{\mathbf{d}}}(SB(\mathcal{A})/X))$ Quotient scheme representing subsystem structures.